### v3

# Equivariant and Invariant Parametrized Topological Complexity

Abstract.
For a $G$-equivariant fibration $p\colon E\to B$, we introduce and study the invariant analogue of Cohen, Farber and Weinberger’s parametrized topological complexity, called the invariant parametrized topological complexity. This notion generalizes the invariant topological complexity introduced by Lubawski and Marzantowicz. We establish the equivariant fibrewise homotopy invariance of this notion and derive several bounds, including a cohomological lower bound and a dimensional upper bound. Additionally, we compare invariant parametrized topological complexity with other well-known invariants. When $G$ is a compact Lie group acting freely on $E$, we show that the invariant parametrized topological complexity of the $G$-fibration $p\colon E\to B$ coincides with the parametrized topological complexity of the induced fibration $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces. Finally, we compute the invariant parametrized topological complexity of equivariant Fadell-Neuwirth fibrations, which measures the complexity of motion planning in presence of obstacles having unknown positions such that the order in which they are placed is irrelevant.
Apart from this, we establish several bounds, including a cohomological lower bound, an equivariant homotopy dimension-connectivity upper bound and various product inequalities for the equivariant sectional category. Applying them, we obtain some interesting results for equivariant and invariant parametrized topological complexity of a $G$-fibration.
Key words and phrases:
equivariant parametrized topological complexity, invariant topological complexity, equivariant sectional category, Lusternik-Schnirelmann category
2020 Mathematics Subject Classification:
55M30, 55R91, 55S40
1. Introduction
The topological complexity of a space $X$, denoted by $\mathrm{TC}(X)$, is defined as the smallest positive integer $k$ such that the product space $X\times X$ can be covered by open sets $\{U_{1},\dots,U_{k}\}$, where each $U_{i}$ admits a continuous section of the free path space fibration
$\pi\colon PX\to X\times X\quad\text{defined by}\quad\pi(\gamma)=(\gamma(0),\gamma(1)),$ (1)
where $PX$ denotes the free path space of $X$ equipped with the compact-open topology. The concept of topological complexity was introduced by Farber in [16] to analyze the computational challenges associated with motion planning algorithms for the configuration space $X$ of a mechanical system. Over the past two decades, this invariant has attracted significant attention and has been a subject of extensive research.
Parameterized motion planning problem
Recently, a novel parametrized approach to the theory of motion planning algorithms was introduced in [8, 9]. This approach provides enhanced universality and flexibility, allowing motion planning algorithms to operate effectively in diverse scenarios by incorporating external conditions. These external conditions are treated as parameters and form an integral part of the algorithm’s input. A parametrized motion planning algorithm takes as input a pair of configurations subject to the same external conditions and produces a continuous motion of the system that remains consistent with these external conditions.
We now briefly define the concept of parametrized topological complexity. For a fibration $p\colon E\to B$, let $E\times_{B}E$ denote the fibre product, which is the space of all pair of points in $E$ that lie in a common fibre of $p$. Let $E^{I}_{B}$ denote the space of all paths in $E$ whose images are contained within in a single fibre. Define the parametrized endpoint map
$\Pi\colon E^{I}_{B}\to E\times_{B}E\quad\text{by}\quad\Pi(\gamma)=(\gamma(0),\gamma(1)).$ (2)
In [8], it is shown that $\Pi$ is a fibration. The parametrized topological complexity of a fibration $p\colon E\to B$, denoted by $\mathrm{TC}[p\colon E\to B]$, is the smallest positive integer $k$ such that there is an open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$, where each $U_{i}$ admits a continuous section of $\Pi$. For further details and interesting computational results for parametrized topological complexity, see [8], [9], [18] and [29]. Additionally, the concept has been extended to fibrewise spaces by García-Calcines in [19]. On the other hand, Crabb [11] established some computational results in the fibrewise setting.
One of the key motivations for introducing this concept was to address the challenge of collision-free motion planning in environments where obstacles have unknown positions in advance. This can be described by the following scenario: A military commander oversees a fleet of $t$ submarines navigating waters with $s$ mines. The positions of these mines change every 24 hours. Each day, the commander must determine a movement plan for each submarine, ensuring that they travel from their current locations to their designated destinations without colliding with either the mines or other submarines. A parametrized motion planning algorithm will take as input the positions of the mines and the current and the desired positions of the submarines and will produced as output a collision-free motion of a fleet. Hence, the complexity of the universal motion planning algorithm in this setting can be described as the parametrized topological complexity of the Fadell-Neuwirth fibration
$p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s),\quad(x_{1},\dots,x_{s},y_{1},\dots,y_{t})\mapsto(x_{1},\dots,x_{s})$
where $F(\mathbb{R}^{d},s)$ is the configuration space of $s$ distinct points lying in $\mathbb{R}^{d}$, see Section 5.
However, in a real-life scenario, the specific order in which the mines are placed should be irrelevant. For the two configurations of mines,
$(x_{1},\dots,x_{s})\quad\text{ and }\quad(x_{\sigma(1)},\dots,x_{\sigma(s)}),$
for any $\sigma$ in the permutation group $\Sigma_{s}$, the military commander should assign the same motion plan for the submarines. This is because both configurations describe the mines being placed at the same set of positions, regardless of their labeling. Thus, we should consider the unordered configuration space $F(\mathbb{R}^{d},s)/\Sigma_{s}$ for the placement of mines. Hence, in this new perspective, the complexity of the universal motion planning algorithm should be described as the parametrized topological complexity of the induced fibration
$\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}$
which is obtained by $p$ by taking the quotient under the natural action of $\Sigma_{s}$ on the configuration spaces. In this paper, we introduce the notion of invariant parametrized topological complexity for a $G$-fibration $p\colon E\to B$, denoted by $\mathrm{TC}^{G}[p\colon E\to B]$, to measure the complexity of parametrized motion planning problem where the order in which the mines are placed is irrelevant.
The invariant parametrized topological complexity is a parametrized analogue of the invariant topological complexity, which was introduced by Lubawski and Marzantowicz [25]. The invariant topological complexity for a $G$-space $X$, denoted by $\mathrm{TC}^{G}(X)$, behaves well with respect to quotients. In particular, if a compact Lie group $G$ acts freely on $X$, then the equality $\mathrm{TC}^{G}(X)=\mathrm{TC}(X/G)$ holds (see [25, Theorem 3.10]). Generalizing this to the parameterized setting we establish the following theorem.
Theorem.
Suppose $G$ is a compact Lie group. Let $p\colon E\to B$ be a $G$-fibration and let $\overline{p}\colon\overline{E}\to\overline{B}$ be the induced fibration between the orbit spaces. If the $G$-action on $E$ is free and $\overline{E}\times\overline{E}$ is hereditary paracompact, then
$\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}].$
Outline of the paper
The aim of this paper is twofold. First, we examine various properties of the equivariant sectional category and equivariant parametrized topological complexity. Using these properties, we develop and analyze the new concept of invariant parametrized topological complexity, which we introduce in Section 4.
In Section 2.1, we study the equivariant sectional category of a $G$-fibration, and establish multiple lower bounds in Theorem 2.2, Proposition 2.4 and Proposition 2.7. We also provide an equivariant homotopy dimension-connectivity upper bound in Theorem 2.12. Afterwards, we establish product inequalities in Proposition 2.14 and Corollary 2.15.
In Section 2.2, we recall the notion of the equivariant LS category of a $G$-space. In this section, we establish a lower bound in terms of fixed point sets, and provide an equivariant homotopy dimension-connectivity upper bound, as stated in Proposition 2.20 and Theorem 2.21, respectively.
Subsequently, Section 2.3 devoted to the equivariant and invariant topological complexity of a $G$-space, and we provide an equivariant homotopy dimension-connectivity upper bound for the former in Theorem 2.23.
In Section 3, we explore various properties of the equivariant parametrized topological complexity of $G$-fibrations $p\colon E\to B$. Our main result Theorem 3.2, characterizes the elements of parametrized motion planning cover as the $G$-compressible subsets of the fibre product $E\times_{B}E$ into the diagonal $\Delta(E)$. Furthermore, we establish some lower bounds and the product inequalities in Proposition 3.4, Theorem 3.6 and Theorem 3.7, respectively.
In Section 4, we introduce the notion of invariant parametrized topological complexity for $G$-fibrations. We establish the fibrewise $G$-homotopy invariance of this notion, and show that it generalizes both the parametrized and invariant topological complexity; see Theorem 4.3 and Proposition 4.4, respectively. For a $G$-fibration $p\colon E\to B$, in Theorem 4.7, we show that the elements of invariant parametrized motion planning cover can be characterized as the $(G\times G)$-compressible subsets of the fibre product $E\times_{B/G}E$ into the saturated diagonal $\mathbb{k}(E)=E\times_{E/G}E$. In Section 4.1, we investigate various properties and bounds for $\mathrm{TC}^{G}[p\colon E\to B]$. For example, we establish inequality under pullbacks (Proposition 4.9), dimensional upper bound (Proposition 4.10), lower bound (Proposition 4.13), cohomological lower bounds (Theorem 4.17 and Theorem 4.19), and product inequality (Theorem 4.20). Finally, we prove one of our main result, Theorem 4.29, which shows that the $\mathrm{TC}^{G}[p\colon E\to B]$ coincides with the parametrized topological complexity of the corresponding orbit fibration, when $G$ acts freely on $E$.
In Section 5, we compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibrations. Specifically, in Theorem 5.6 and Theorem 5.11, we establish:
Theorem.
Suppose $s\geq 2$, $t\geq 1$ and $d\geq 3$. Then
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\begin{cases}2t+s,&\text{if $d$ is odd},\\ \text{either }2t+s-1\text{ or }2t+s&\text{if $d$ is even}.\end{cases}$
Theorem.
Suppose $s\geq 2$ and $t\geq 2$. Then
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=2t+s-1.$
2. Preliminaries
In this section, we systematically introduce and study various numerical invariants: equivariant sectional category, equivariant LS-category, equivariant topological complexity, $A$-Lusternik-Schnirelmann $G$-category, and invariant topological complexity.
2.1. Equivariant sectional category
Schwarz [32] introduced and studied the notion of sectional category of a fibration, and later by Bernstein and Ganea in [3] for any map. The corresponding equivariant analogue was introduced by Colman and Grant in [10].
Definition 2.1 ([10, Definition 4.1]).
Let $p\colon E\to B$ be a $G$-map. The equivariant sectional category of $p$, denoted by $\mathrm{secat}_{G}(p)$, is the least positive integer $k$ such that there is a $G$-invariant open cover $\{U_{1},\ldots,U_{k}\}$ of $B$ and $G$-maps $s_{i}\colon U_{i}\to E$, for $i=1,\ldots,k$, such that $p\circ s_{i}\simeq_{G}i_{U_{i}}$, where $i_{U_{i}}\colon U_{i}\hookrightarrow B$ is the inclusion map.
First we establish a cohomological lower bound on the equivariant sectional category of a $G$-map using Borel cohomology. To the best of our knowledge, such a bound has not been documented in the literature. We believe that this result must already be known to experts in the field. Nevertheless, we provide a thorough proof of this result here.
Suppose $EG\to BG$ is a universal principal $G$-bundle. For a $G$-space $X$, let $X^{h}_{G}$ be the homotopy orbit space of $X$ defined as
$X^{h}_{G}:=EG\times_{G}X=\frac{EG\times X}{(eg,x)\sim(e,g^{-1}x)},\quad\text{for }e\in EG,g\in G,x\in X$
and the Borel $G$-equivariant cohomology $H^{*}_{G}(X;R)$ of $X$ with coefficients in a commutative ring $R$ is defined as $H^{*}_{G}(X;R):=H^{*}(X^{h}_{G};R)$. We note that for a $G$-map $p\colon E\to B$, there is an induced map $p^{h}_{G}\colon E^{h}_{G}\to B^{h}_{G}$.
Theorem 2.2 (Cohomological lower bound).
Suppose $p\colon E\to B$ is a $G$-map. If there are cohomology classes $u_{1},\dots,u_{k}\in\widetilde{H}^{*}_{G}(B;R)$ (for any commutative ring $R$) with
$(p^{h}_{G})^{*}(u_{1})=\cdots=(p^{h}_{G})^{*}(u_{k})=0\quad\text{and}\quad u_{1}\smile\dots\smile u_{k}\neq 0,$
then $\mathrm{secat}_{G}(p)>k$.
Proof.
Suppose $\mathrm{secat}_{G}(p)\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $B$ such that each $U_{i}$ admits a $G$-equivariant homotopy section $s_{i}$ of $p$. Let $j_{i}\colon U_{i}\hookrightarrow B$ be the inclusion map. Then
$((j_{i})^{h}_{G})^{*}(u_{i})=((s_{i})^{h}_{G})^{*}\left((p^{h}_{G})^{*}(u_{i})\right)=0$
since $p\circ s_{i}\simeq_{G}j_{i}$ implies $((j_{i})^{h}_{G})^{*}=((s_{i})^{h}_{G})^{*}\circ(p^{h}_{G})^{*}$. Hence, the long exact sequence in cohomology associated to the pair $(B^{h}_{G},(U_{i})^{h}_{G})$ gives an element $v_{i}\in H^{*}(B^{h}_{G},(U_{i})^{h}_{G};R)$ such that $((q_{i})^{h}_{G})^{*}(v_{i})=u_{i}$, where $q_{i}\colon B\hookrightarrow(B,U_{i})$ is the inclusion map. Hence, we get
$v_{1}\smile\cdots\smile v_{k}\in H^{*}(B^{h}_{G},\cup_{i=1}(U_{i})^{h}_{G};R)=H^{*}(B^{h}_{G},B^{h}_{G};R)=0.$
Moreover, by the naturality of cup products, we have $(q^{h}_{G})^{*}(v_{1}\smile\cdots\smile v_{k})=u_{1}\smile\cdots\smile u_{k}$, where $q\colon B\hookrightarrow(B,B)$ is the inclusion map. Hence, $u_{1}\smile\cdots\smile u_{k}=0$. ∎
Remark 2.3.
  1. (1)
    Observe that if $G$ acts trivially on $X$, then the lower bound in Theorem 2.2 recovers the cohomological lower bound given by Schawrz in [32, Theorem 4].
  2. (2)
    Note the following commutative diagram of $G$-maps
    ${X}$${PX}$${X\times X,}$$\scriptstyle{\Delta}$$\scriptstyle{h}$$\scriptstyle{\pi}$
    where $h$ is a $G$-homotopy equivalence. Then, the lower bound in Theorem 2.2 recovers the bound [10, Theorem 5.15] on $\mathrm{TC}_{G}(X)$, obtained by Colman and Grant. More generally, it also recovers the bound [12, Theorem 4.25] on the equivariant parametrized topological complexity, obtained by the second author.
In practice, however, the difficulty of computing cup products in Borel cohomology (or more generally, in any equivariant cohomology) makes the problem cumbersome. We can then ask whether non-equivariant cohomological bounds can be utilized in some way. When $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration, we will show that the sectional category of $p$ and the sectional category of the induced fibration $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces are lower bounds for the equivariant sectional category of $p$. Note that $\overline{p}$ fits into the commutative diagram
[…]
where $\pi_{B}\colon B\to\overline{B}$ and $\pi_{E}\colon E\to\overline{E}$ are orbit maps.
Proposition 2.4.
Let $p\colon E\to B$ be a $G$-fibration. Then $\mathrm{secat}(\overline{p})\leq\mathrm{secat}_{G}(p).$
Proof.
Suppose $U$ is a $G$-invariant open subset of $B$ with a $G$-equivariant section $s$ of $p$ over $U$. As the orbit map $\pi_{B}\colon B\to\overline{B}$ is open, we have $\overline{U}:=\pi_{B}(U)$ is an open subset of $\overline{B}$. As $U$ is $G$-invariant, it follows $U$ is saturated with respect to $\pi_{B}$. Hence, $\pi_{B}\colon U\to\overline{U}$ is a quotient map. Then, by universal property of quotient maps, there exists a unique continuous map $\overline{s}\colon\overline{U}\to\overline{E}$ such that the following diagram
[…]
commutes. Then
$\overline{p}(\overline{s}(\overline{b}))=\overline{p}(\overline{s}(\pi_{B}(b)))=\overline{p}(\pi_{E}(s(b)))=\pi_{B}(p(s(b)))=\pi_{B}(b)=\overline{b}$
implies $\overline{s}$ is a section of $\overline{p}$ over $\overline{U}$. Thus, the result follows since $\pi_{B}\colon B\to\overline{B}$ is surjective. ∎
Theorem 2.5.
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration. Then $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration. Furthermore, if $E$ and $B$ are Hausdorff, and $G$ acts freely on $B$, then
$\mathrm{secat}(\overline{p})=\mathrm{secat}_{G}(p).$
Proof.
By [20, Corollary 2], it follows that $\overline{p}$ is a fibration. Note that the inequality $\mathrm{secat}(\overline{p})\leq\mathrm{secat}_{G}(p)$ follows from Proposition 2.4. Now we will show the reverse inequality. Note that $G$ also acts freely on $E$ since $p$ is a $G$-map. Hence, the diagram (3) is a pullback in the category of $G$-spaces, see the proof of [7, Theorem II.7.3]. Suppose $\overline{U}$ is an open subset of $\overline{B}$ and $\overline{s}\colon\overline{U}\to\overline{E}$ is a section of $\overline{p}$. Let $U=\pi_{B}^{-1}(\overline{U})$. Then, by the universal property of pullbacks, there exists a unique $G$-map $s\colon U\to E$ such that $\pi_{E}\circ s=\overline{s}\circ\pi_{B}\colon U\to\overline{E}$ and $p\circ s=i_{U}\colon U\to B$. Hence, $\mathrm{secat}_{G}(p)\leq\mathrm{secat}(\overline{p})$. ∎
Remark 2.6.
For any space $X$, the free path space fibration $\pi\colon PX\to X\times X$ is a $\mathbb{Z}_{2}$-fibration, with $\mathbb{Z}_{2}$-action on $PX$ given by reversal of paths and on $X\times X$ by transposition of factors, see [21, Example 2.6]. Hence, by Theorem 2.5, we get $\mathrm{secat}(\overline{\pi})\leq\mathrm{secat}_{\mathbb{Z}_{2}}(\pi)$. Thus, we recover the cohomological lower bound on the symmetrized topological complexity $\mathrm{TC}^{\Sigma}(X)$ in [21, Theorem 4.6] since the following commutative diagram
[…]
implies the nilpotency of the kernel of $\overline{\Delta}$ and $\overline{\pi}$ are the same.
Suppose $X$ is a $G$-space. For a subgroup $H$ of $G$, define the $H$-invariant subspace of $X$ as
$X^{H}:=\{x\in X\mid h\cdot x=x\text{ for all }h\in H\}.$
Proposition 2.7.
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-map, then
$\mathrm{secat}_{K}(p^{H})\leq\mathrm{secat}_{G}(p).$
Proof.
Suppose $s\colon U\to E$ is a $G$-equivariant section of $p$. Define $V=U\cap E^{H}$. Note that $V$ is $H$-fixed points of $U$ and is $K$-invariant. Since $s$ is $G$-equivariant, it takes $H$-fixed points to $H$-fixed points, and hence restricts to a $K$-equivariant map $\left.s\right|_{V}\colon V\to B^{H}$. It is clear that $\left.s\right|_{V}$ is a section of $p^{H}$. ∎
Corollary 2.8.
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration. Then
  1. (1)
    the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a fibration for all closed subgroups $H$ of $G$, and
    $\mathrm{secat}(p^{H})\leq\mathrm{secat}_{G}(p).$
  2. (2)
    $p\colon E\to B$ is a $K$-fibration for all closed subgroups $K$ of $G$, and
    $\mathrm{secat}(p)\leq\mathrm{secat}_{K}(p)\leq\mathrm{secat}_{G}(p).$
Proof.
Suppose $T$ is the trivial subgroup of $G$. By [20, Theorem 4], we know that for a $G$-fibration $p$, the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is fibration for all closed subgroups $H$ of $G$. Hence, by taking $K=T$ in Proposition 2.7, it follows that $\mathrm{secat}(p^{H})=\mathrm{secat}_{T}(p^{H})\leq\mathrm{secat}_{G}(p)$.
By [20, Theorem 3], we know that $p$ is a $K$-fibration for all closed subgroups $K$ of $G$. Hence, by taking $H=T$ in Proposition 2.7, we get the subgroup inequality $\mathrm{secat}_{K}(p)=\mathrm{secat}_{K}(p^{T})\leq\mathrm{secat}_{G}(p)$. Note that $T$ is a closed subgroup of compact Hausdorff topological group $K$. Hence, applying the subgroup inequality for the $K$-fibration $p$, we get $\mathrm{secat}(p)=\mathrm{secat}_{T}(p)\leq\mathrm{secat}_{K}(p)$. ∎
The following proposition states some basic properties of the equivariant sectional category. Proofs are left to the reader. For analogous results concerning the non-equivariant sectional category, refer to [22, Lemma 2.1].
Proposition 2.9.
Suppose $p\colon E\to B$ is a $G$-map.
  1. (1)
    If $p^{\prime}\colon E\to B$ is $G$-homotopic to $p$, then $\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(p)$.
  2. (2)
    If $h\colon E^{\prime}\to E$ is $G$-homotopy equivalence, then $\mathrm{secat}_{G}(p\circ h)=\mathrm{secat}_{G}(p)$.
  3. (3)
    If $f\colon B\to B^{\prime}$ is a $G$-homotopy equivalence, then $\mathrm{secat}_{G}(f\circ p)=\mathrm{secat}_{G}(p)$.
Corollary 2.10.
Suppose $p\colon E\to B$ is a $G$-fibration. If $g\colon B^{\prime}\to B$ is a $G$-homotopy equivalence and $p^{\prime}\colon E^{\prime}\to B^{\prime}$ is the pullback of $p$ along $g$, then
$\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(p).$
Proof.
Suppose the following diagram is a pullback
[…]
Since $g$ is a $G$-homotopy equivalence and $p$ is a $G$-fibration, it follows that $h$ is also a $G$-homotopy equivalence. Hence, we get
$\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(g\circ p^{\prime})=\mathrm{secat}_{G}(p\circ h)=\mathrm{secat}_{G}(p).$
by Proposition 2.9. ∎
Generalizing Schwarz’s dimension-connectivity upper bound on the sectional category, Grant established the corresponding equivariant analogue for the equivariant sectional category in [21, Theorem 3.5]. We extend this approach to derive an equivariant homotopy dimension-connectivity upper bound for equivariant sectional category. To achieve this, we first introduce the notion of $G$-homotopy dimension for $G$-$\mathrm{CW}$-complexes.
Definition 2.11.
Suppose $X$ is a $G$-$\mathrm{CW}$-complex. The $G$-homotopy dimension of $X$, denoted $\mathrm{hdim}_{G}(X)$, is defined to be
$\mathrm{hdim}_{G}(X):=\min\{\dim(X^{\prime})\mid X^{\prime}\text{ is a $G$-}\mathrm{CW}\text{-complex},X^{\prime}\simeq_{G}X\}.$
Theorem 2.12.
Suppose $p\colon E\to B$ is a Serre $G$-fibration with fibre $F$, whose base $B$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$. If there exists $s\geq 0$ such that the fibre of $p^{H}\colon E^{H}\to B^{H}$ is $(s-1)$-connected for all subgroups $H$ of $G$, then
$\mathrm{secat}_{G}(p)<\frac{\mathrm{hdim_{G}(B)+1}}{s+1}+1.$
Proof.
It is enough to show that for any $G$-$\mathrm{CW}$-complex $B^{\prime}$ which is $G$-homotopy equivalent to $B$, we have
$\mathrm{secat}_{G}(p)<\frac{\mathrm{dim(B^{\prime})+1}}{s+1}+1.$
Suppose $f\colon B^{\prime}\to B$ is a $G$-homotopy between $G$-$\mathrm{CW}$-complexes $B^{\prime}$ and $B$, and $p^{\prime}\colon E^{\prime}\to B^{\prime}$ is the pullback of $p$ along $f$. Then, by Corollary 2.10, we have $\mathrm{secat}_{G}(p^{\prime})=\mathrm{secat}_{G}(p)$. Since the fibre of $p^{\prime}$ is also $F$, we get
$\mathrm{secat}_{G}(p^{\prime})<\frac{\mathrm{dim(B^{\prime})+1}}{s+1}+1,$
by [21, Theorem 3.5]. ∎
Our next aim is to establish product inequalities for the equivariant sectional category.
Definition 2.13.
A $G$-space $X$ is called $G$-completely normal if for any two $G$-invariant subsets $A$ and $B$ of $X$ with $\overline{A}\cap B=A\cap\overline{B}=\emptyset$, there exist disjoint $G$-invariant open subsets of $X$ containing $A$ and $B$, respectively.
Proposition 2.14.
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is a compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts on $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$ diagonally. Furthermore, if $B_{1}$ and $B_{2}$ are Hausdorff, and $B_{1}\times B_{2}$ is completely normal, then
$\mathrm{secat}_{G}(p_{1}\times p_{2})\leq\mathrm{secat}_{G}(p_{1})+\mathrm{secat}_{G}(p_{2})-1.$
Proof.
Suppose $G$ is compact Hausdorff. Then identifying $G$ with the diagonal subgroup of $G\times G$, we see that it is a closed subgroup of $G\times G$. Hence, by [20, Theorem 3], it follows that $p_{1}\times p_{2}\colon E_{2}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$.
If $B_{1}$ and $B_{2}$ are Hausdorff, and $B_{1}\times B_{2}$ is completely normal, then [10, Lemma 3.12] implies $B_{1}\times B_{2}$ is $(G\times G)$-completely normal. Hence, the desired inequality
$\mathrm{secat}_{G}(p_{1}\times p_{2})\leq\mathrm{secat}_{G\times G}(p_{1}\times p_{2})\leq\mathrm{secat}_{G}(p_{1})+\mathrm{secat}_{G}(p_{2})-1$
follows from Corollary 2.8 (2) and [1, Proposition 3.7]. ∎
Corollary 2.15.
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $B$ is Hausdorff and $B\times B$ is completely normal, then
$\mathrm{secat}_{G}(p)\leq\mathrm{secat}_{G}(p_{1})+\mathrm{secat}_{G}(p_{2})-1.$
Proof.
Note that the following diagram
[…]
is a pullback in the category of $G$-spaces, where $\Delta\colon B\to B\times B$ is the diagonal map. In Proposition 2.14, we showed that $p_{1}\times p_{2}$ is a $G$-fibration if $G$ is compact Hausdorff. Hence, $p$ is a $G$-fibration. Thus, the desired inequality
$\mathrm{secat}_{G}(p)\leq\mathrm{secat}_{G}(p_{1}\times p_{2})\leq\mathrm{secat}_{G}(p_{1})+\mathrm{secat}_{G}(p_{2})-1$
follows from [10, Proposition 4.3] and Proposition 2.14. ∎
2.2. Equivariant LS-category
The notion of Lusternik-Schnirelmann (LS) category was introduced by Lusternik and Schnirelmann in [26]. In this section, we present the corresponding equivariant analogue.
Definition 2.16.
A $G$-invariant subset $U$ of a $G$-space $X$ is said to be $G$-categorical if the inclusion map $i_{U}\colon U\hookrightarrow X$ is $G$-homotopy equivalent to a map which takes values in a single orbit.
Definition 2.17.
[14] The equivariant LS-category of a $G$-space $X$, denoted by $\mathrm{cat}_{G}(X)$, is the least positive integer $k$ such that there exists a $G$-categorical open cover $\{U_{1},\dots,U_{k}\}$ of $X$.
Definition 2.18.
A $G$-space $X$ is said to be $G$-connected if $X^{H}$ is path-connected for every closed subgroup $H$ of $G$.
Let $X$ be a $G$-space, and $x_{0}\in X$. Define the path space of $(X,x_{0})$ as
$P_{x_{0}}X=\{\alpha\colon I\to X\mid\alpha(0)=x_{0}\}.$
Then the map $e_{X}\colon P_{x_{0}}X\to X$, given by $e_{X}(\alpha)=\alpha(1)$, is a fibration. Moreover, if the point $x_{0}$ is fixed under the $G$-action, then $e_{X}$ is a $G$-fibration, where $P_{x_{0}}X$ admits a $G$-action via $(g\cdot\alpha)(t):=g\cdot\alpha(t)$. We note that the fibre of $e_{X}$ is the based loop space $\Omega X=(e_{X})^{-1}(x_{0})$ of $X$, and the $G$-action on $P_{x_{0}}X$ restricts to a $G$-action on $\Omega X$. Furthermore, we have a commutative diagram of $G$-maps
${\{x_{0}\}}$${P_{x_{0}}X}$${X,}$$\scriptstyle{h}$$\scriptstyle{i}$$\scriptstyle{e_{X}}$ (4)
where $h$ is a $G$-homotopy equivalence and $i\colon\{x_{0}\}\hookrightarrow X$ is the inclusion map.
Lemma 2.19 ([10, Corollary 4.7]).
If $X$ is a $G$-space such that $X$ is $G$-connected and $x_{0}\in X^{G}$, then $\mathrm{cat}_{G}(X)=\mathrm{secat}_{G}(e_{X}).$
We now present inequalities relating $\mathrm{cat}_{G}(X)$ to the non-equivariant category of fixed point sets and to the equivariant category of $X$ viewed as a $K$-space, for each closed subgroup $K$ of $G$.
Proposition 2.20.
Suppose $X$ is a $G$-connected space with $X^{G}\neq\emptyset$. If $H$ and $K$ are closed subgroups of $X$ such that $X^{H}$ is $K$-invariant, then
$\mathrm{cat}_{K}(X^{H})\leq\mathrm{cat}_{G}(X).$
In particular, if $G$ is Hausdorff, then
  1. (1)
    $\mathrm{cat}(X^{H})\leq\mathrm{cat}_{G}(X)$ for all closed subgroups $H$ of $G$.
  2. (2)
    $\mathrm{cat}_{K}(X)\leq\mathrm{cat}_{G}(X)$ for all closed subgroups $K$ of $G$.
Proof.
We note that $X^{H}$ is $K$-connected since $X$ is $G$-connected, and $H$ and $K$ are closed subgroups of $G$. If $x_{0}\in X^{G}$, then $x_{0}\in(X^{H})^{K}=X^{H\cap K}$. Hence, by Lemma 2.19, it is enough to show that $\mathrm{secat}_{K}(e_{X^{H}})\leq\mathrm{secat}_{G}(e_{X})$.
Suppose $U$ is a $G$-invariant open subset of $X$ and $s\colon U\to P_{x_{0}}X$ is $G$-equivariant section of $e_{X}$. Set $V:=U\cap X^{H}$. Then $V$ is a $K$-invariant open subset of $X^{H}$. As $s$ is $G$-equivariant, it restricts to a $K$-equivariant map $\left.s\right|_{V}\colon V\to(P_{x_{0}}X)^{H}=P_{x_{0}}(X^{H})$. Clearly, $\left.s\right|_{V}$ is a $K$-equivariant section of $e_{X^{H}}\colon P_{x_{0}}(X^{H})\to X^{H}$. ∎
Now, as a consequence of Theorem 2.12, we obtain an equivariant homotopy dimension-connectivity upper bound for equivariant LS category.
Theorem 2.21.
Suppose $X$ is a $G$-$\mathrm{CW}$-complex of dimension at least $2$ such that $X^{G}\neq\emptyset$. If there exists $s\geq 0$ such that $X^{H}$ is $s$-connected for all subgroups $H$ of $G$, then
$\mathrm{cat}_{G}(X)<\frac{\mathrm{hdim}_{G}(X)+1}{s+1}+1.$
Proof.
If $x_{0}\in X^{G}$, then $e_{X}\colon P_{x_{0}}X\to X$ is a $G$-fibration with fibre $\Omega X$ which also admits a $G$-action. Note that $(\Omega X)^{H}=\Omega(X^{H})$. Since $X^{H}$ is $s$-connected, the loop space $\Omega(X^{H})$ is $(s-1)$-connected. Hence, by Theorem 2.12, we get
$\mathrm{secat}_{G}(e_{X})<\frac{\mathrm{hdim}_{G}(X)+1}{s+1}+1.$
As $X^{H}$ is $s$-connected, it follows that $X^{H}$ is path-connected. Hence, $X$ is $G$-connected, and the theorem follows by Lemma 2.19. ∎
2.3. Equivariant and invariant topological complexity
We recall the concept of equivariant topological complexity introduced by Colman and Grant in [10]. Let $X$ be a $G$-space. Observe that the free path space $PX$ admits a $G$-action via $(g\cdot\alpha)(t):=g\cdot\alpha(t)$. Similarly, the product space $X^{k}$ is a $G$-space with the diagonal action. The fibration
$e_{k,X}\colon PX\to X^{k},\quad\alpha\mapsto\left(\alpha(0),\alpha\left(\frac{1}{k-1}\right),\ldots,\alpha\left(\frac{i}{k-1}\right),\ldots,\alpha\left(\frac{k-2}{k-1}\right),\alpha(1)\right)$
is a $G$-fibration.
Definition 2.22.
The sequential equivariant topological complexity of a $G$-space $X$ is defined as
$\mathrm{TC}_{k,G}(X):=\mathrm{secat}_{G}(e_{k,X}).$
In particular, when $k=2$, we will denote $e_{2,X}$ by $\pi$ and $\mathrm{TC}_{2,G}(X)$ by $\mathrm{TC}_{G}(X)$.
In [1, Proposition 3.40], Sarkar and the authors of this paper provided a dimension-connectivity upper bound on the sequential equivariant topological complexity. We improve their result by establishing an equivariant homotopy dimension-connectivity upper bound. We omit the proof, as it is similar to the original and follows from the homotopy dimension-connectivity upper bound on the equivariant sectional category in Theorem 2.12.
Theorem 2.23.
Let $X$ be a $G$-$\mathrm{CW}$-complex of dimension atleast 1 such that $X^{H}$ is $m$-connected for all subgroups $H\leq G$. Then
$\mathrm{TC}_{k,G}(X)<\frac{k\ \mathrm{hdim}_{G}(X)+1}{m+1}+1.$
It is important to note that the equivariant topological complexity of $G$-spaces does not necessarily relate to the topological complexity of their orbit spaces. However, Lubawski and Marzantowicz provided an alternative definition of equivariant topological complexity, designed to facilitate such a comparison. We now present their definition and recall the corresponding result.
Suppose $X$ is a $G$-space. Let $\pi_{X}:X\to X/G$ denote the orbit map.
$PX\times_{X/G}PX:=\{(\gamma,\delta)\in PX\times PX\mid G\cdot\gamma(1)=G\cdot\delta(0)\}$
That is the following diagram
${PX\times_{X/G}PX}$${PX}$${PX}$${X/G}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{X}\circ e_{0}}$$\scriptstyle{\pi_{X}\circ e_{1}}$
is a pullback. Define the map
$\mathfrak{p}\colon PX\times_{X/G}PX\to X\times X,\quad(\gamma,\delta)\mapsto(\gamma(0),\delta(1)).$
It was shown in [25, Proposition 3.7] that the map $\mathfrak{p}$ is a $(G\times G)$-fibration.
Definition 2.24.
Let $X$ be a $G$-space. The invariant topological complexity of $X$ denoted by $\mathrm{TC}^{G}(X)$, is defined as
$\mathrm{TC}^{G}(X):=\mathrm{secat}_{G\times G}(\mathfrak{p}).$
The following theorem relates the invariant topological complexity of a free $G$-space $X$ with that of the topological complexity of its corresponding orbit space.
Theorem 2.25 ([25, Theorem 3.9 and 3.10]).
Let $G$ be a compact Lie group and $X$ be a compact $G$-ANR. Then
$\mathrm{TC}(X/G)\leq\mathrm{TC}^{G}(X).$
Moreover, if $X$ has one orbit type, then
$\mathrm{TC}^{G}(X)=\mathrm{TC}(X/G).$
2.4. Clapp-Puppe invariant of Lusternik-Schnirelmann type
Definition 2.26.
Let $A$ be a $G$-invariant closed subset of a $G$-space $X$. A $G$-invariant open subset of $X$ is said to be $G$-compressible into $A$ if the inclusion map $i_{U}\colon U\rightarrow X$ is $G$-homotopic to a $G$-map $c\colon U\to X$ which takes values in $A$.
Definition 2.27.
Let $A$ be a $G$-invariant closed subset of a $G$-space $X$. The $A$-Lusternik-Schnirelmann $G$-category of $X$, denoted ${}_{A}\mathrm{cat}_{G}(X)$, is the least positive integer $k$ such that there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $X$ such that each $U_{i}$ is $G$-compressible into $A$.
Colman and Grant in [10, Lemma 5.14] showed that for a $G$-invariant open subset $U$ of $X\times X$ the following are equivalent:
  1. (1)
    there exists a $G$-equivariant section of $e_{X}\colon PX\to X\times X$ over $U$,
  2. (2)
    $U$ is $G$-compressible into the diagonal $\Delta(X)\subset X\times X$.
In particular,
$\mathrm{TC}_{G}(X)=~{}_{\Delta(X)}\mathrm{cat}_{G}(X\times X).$
Later, Lubawski and Marzantowicz in [25, Lemma 3.8] showed a similar result for invariant topological complexity. More precisely, for a $(G\times G)$-invariant open subset of $U$ of $X\times X$ the following are equivalent:
  1. (1)
    there exists a $(G\times G)$-equivariant section of $\mathfrak{p}\colon PX\times_{X/G}X\to X\times X$ over $U$,
  2. (2)
    $U$ is $(G\times G)$-compressible into the saturation of the diagonal $\mathbb{k}(X):=(G\times G)\cdot\Delta(X)\subset X\times X$.
In particular,
$\mathrm{TC}^{G}(X)=~{}_{\mathbb{k}(X)}\mathrm{cat}_{G\times G}(X\times X).$
In Section 3 and Section 4, we give analogous results for equivariant parametrized topological complexity and invariant parametrized topological complexity, respectively. We use these results to prove Theorem 4.29.
3. Equivariant parametrized topological complexity
For a $G$-fibration $p\colon E\to B$, consider the subspace $E^{I}_{B}$ of the free path space $E^{I}$ of $E$ defined by
$E^{I}_{B}:=\{\gamma\in E^{I}\mid\gamma(t)\in p^{-1}(b)~{}\text{for some}~{}b\in B~{}\text{and for all}~{}t\in[0,1]\}.$
Consider the pullback corresponding to the fibration $p\colon E\to B$ defined by
$E\times_{B}E=\{(e_{1},e_{2})\in E\times E\mid p(e_{1})=p(e_{2})\}.$
It is clear that the $G$-action on $E^{I}$ given by
$(g\cdot\gamma)(t):=g\cdot\gamma(t)\quad\text{for all }g\in G,\gamma\in E^{I},t\in I;$
and the diagonal action of $G$ on $E\times E$ restricts to $E^{I}_{B}$ and $E\times_{B}E$, respectively. Then the map
$\Pi\colon E^{I}_{B}\to E\times_{B}E,\quad\Pi(\gamma)=(\gamma(0),\gamma(1))$ (5)
is a $G$-fibration, see [12, Corollary 4.3].
Definition 3.1 ([12, Definition 4.1]).
The equivariant parametrized topological complexity of a $G$-fibration $p\colon E\to B$, denoted by $\mathrm{TC}_{G}[p\colon E\to B]$, is defined as
$\mathrm{TC}_{G}[p\colon E\to B]:=\mathrm{secat}_{G}\left(\Pi\right).$
Suppose $\Delta\colon E\to E\times E$ is the diagonal map. Then it is clear that the image $\Delta(E)$ is a $G$-invariant subset of $E\times_{B}E$. In the next theorem, we prove the parametrized analogue of [10, Lemma 5.14] in the equivariant setting.
Theorem 3.2.
Let $p\colon E\to B$ be a $G$-fibration. For a $G$-invariant (not necessarily open) subset $U$ of $E\times_{B}E$ the following are equivalent:
  1. (1)
    there exists a $G$-equivariant section of $\Pi\colon E^{I}_{B}\to E\times_{B}E$ over $U$.
  2. (2)
    there exists a $G$-homotopy between the inclusion map $i_{U}\colon U\hookrightarrow E\times_{B}E$ and a $G$-map $f\colon U\to E\times_{B}E$ which takes values in $\Delta(E)$.
Proof.
$(1)\implies(2)$. Suppose $s\colon U\to E^{I}_{B}$ is a $G$-equivariant section of $\Pi$. Let $H\colon E^{I}_{B}\times I\to E^{I}_{B}$ be given by
$H(\gamma,t)(s)=\gamma(s(1-t)),\quad\text{for }\gamma\in E^{I}_{B}\text{ and }s,t\in I.$
It is clear that $H(\gamma,t)\in E^{I}_{B}$ for all $\gamma\in E^{I}_{B}$ and $t\in I$. Hence, $H$ is well-defined. Clearly, $H$ is $G$-equivariant such that $H(\gamma,0)=\gamma$ and $H(\gamma,1)=c_{\gamma(0)}$, where $c_{e}$ is the constant path in $E$ taking the value $e\in E$. Then
$F:=\Pi\circ H\circ(s\times\mathrm{id}_{I})\colon U\times I\to E\times_{B}E$
is a $G$-homotopy such that $F_{0}=\Pi\circ\mathrm{id}_{E^{I}_{B}}\circ s=i_{U}$ and $F_{1}(u)=\Pi(H_{1}(s(u)))=\Pi(c_{s(u)(0)})=(s(u)(0),s(u)(0))\in\Delta(E)$. Hence, $F_{1}$ is the desired map.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B}E$ is a $G$-homotopy between $f$ and $i_{U}$. Let $s\colon U\to E^{I}_{B}$ be the $G$-map given by $s(u)=c_{\pi_{1}(f(u))}=c_{\pi_{2}(f(u))}$, where $\pi_{i}\colon E\times E\to E$ is the projection map onto the $i$-th factor. By $G$-homotopy lifting property of $\Pi$, there exists a $G$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}$ such that the following diagram
[…]
commutes. Then $\Pi\circ\widetilde{H}_{1}=H_{1}=i_{U}$ implies $\widetilde{H}_{1}$ is a $G$-equivariant section of $\Pi$ over $U$. ∎
As a consequence to the previous theorem we can now express the equivariant parametrized topological complexity as the equivariant $\Delta(E)$-LS category of the fibre product.
Corollary 3.3.
For a $G$-fibration $p\colon E\to B$, we have
$\mathrm{TC}_{G}[p\colon E\to B]=~{}_{\Delta(E)}\mathrm{cat}_{G}(E\times_{B}E).$
Proposition 3.4.
Suppose $p\colon E\to B$ is a $G$-fibration. If $H$ and $K$ are subgroups of $G$ such that the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a $K$-fibration, then
$\mathrm{TC}_{K}[p^{H}\colon E^{H}\to B^{H}]\leq\mathrm{TC}_{G}[p\colon E\to B].$
Proof.
Suppose $\Pi\colon E^{I}_{B}\to E\times_{B}E$ is $G$-equivariant parametrized fibration corresponding to $p$. Then it is easily checked that
$(E^{I}_{B})^{H}=(E^{H})^{I}_{B^{H}}\quad\text{and}\quad(E\times_{B}E)^{H}=E^{H}\times_{B^{H}}E^{H},$
and the $K$-equivariant parameterized fibration corresponding to $p^{H}$ is given by $\Pi^{H}$. Hence, it follows that
$\mathrm{TC}_{K}[p^{H}\colon E^{H}\to B^{H}]=\mathrm{secat}_{K}(\Pi^{H})\leq\mathrm{secat}_{G}(\Pi)=\mathrm{TC}_{G}[p\colon E\to B]$
by Proposition 2.7. ∎
Applying Proposition 3.4 and Corollary 2.8, we obtain the following corollary.
Corollary 3.5.
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration. Then
  1. (1)
    the fixed point map $p^{H}\colon E^{H}\to B^{H}$ is a fibration for all closed subgroups $H$ of $G$, and
    $\mathrm{TC}[p^{H}\colon E^{H}\to B^{H}]\leq\mathrm{TC}_{G}[p\colon E\to B].$
  2. (2)
    $p\colon E\to B$ is a $K$-fibration for all closed subgroups $K$ of $G$, and
    $\mathrm{TC}[p\colon E\to B]\leq\mathrm{TC}_{H}[p\colon E\to B]\leq\mathrm{TC}_{G}[p\colon E\to B].$
A cohomological lower bound for the equivariant parametrized topological complexity was established by the second author in [12, Theorem 4.5] using Borel cohomology. In the following theorem, we provide an alternative cohomological lower bound that is easier to compute since it relies on non-equivariant cohomology.
Let $E_{B,G}:=(E\times_{B}E)/G$ and let $d_{G}E\subseteq E_{B,G}$ denote the image of the diagonal subspace $\Delta(E)\subseteq E\times_{B}E$ under the orbit map $\rho\colon E\times_{B}E\to E_{B,G}$.
Theorem 3.6.
Suppose $p\colon E\to B$ is a $G$-fibration. If there exists cohomology classes $u_{1},\dots,u_{k}\in H^{*}(E_{B,G};R)$ (for any commutative ring $R$) such that
  1. (1)
    $u_{i}$ restricts to zero in $H^{*}(d_{G}E;R)$ for $i=1,\dots,k$;
  2. (2)
    $u_{1}\smile\dots\smile u_{k}\neq 0$ in $H^{*}(E_{B,G};R)$,
then $\mathrm{TC}_{G}[p\colon E\to B]>k$.
Proof.
Suppose $\mathrm{TC}_{G}[p\colon E\to B]\leq k$. Then there exists a $G$-invariant open cover $\{U_{1},\dots,U_{k}\}$ of $E\times_{B}E$ such that each $U_{i}$ admits a $G$-equivariant section of $\Pi$. By Theorem 3.2, for each $i=1,\dots,k$, there exists a $G$-homotopy $H_{i}\colon U_{i}\times I\to E\times_{B}E$ from the inclusion map $j_{U_{i}}\colon U_{i}\hookrightarrow E\times_{B}E$ to a $G$-map $f_{i}\colon U_{i}\to E\times_{B}E$ which takes values in $\Delta(E)$. Let $\overline{U_{i}}:=\rho(U_{i})$. As $I$ is locally compact, $H_{i}$ induces a homotopy $\overline{H_{i}}\colon\overline{U_{i}}\times I\to E_{B,G}$ from the inclusion map $j_{\overline{U_{i}}}\colon\overline{U_{i}}\hookrightarrow E_{B,G}$ to a map $\overline{f_{i}}\colon\overline{U_{i}}\to E_{B,G}$ which takes values in $d_{G}E$. Thus, the following diagram
[…]
is commutative. Hence, by hypothesis (1), each $u_{i}$ restricts to zero in $H^{*}(\overline{U_{i}};R)$. By long exact sequence of the pair $(E_{B,G},d_{G}E)$, there exists classes $v_{i}\in H^{*}(E_{B,G},\overline{U_{i}};R)$ such that $v_{i}$ maps to $u_{i}$ under the coboundary map $H^{*}(E_{B,G},\overline{U_{i}};R)\to H^{*}(E_{B,G};R)$. Hence, we get
$v_{1}\smile\dots\smile v_{k}\in H^{*}(E_{B,G},\cup_{i=1}^{k}\overline{U_{i}};R)=H^{*}(E_{B,G},E_{B,G};R)=0.$
Thus, by the naturality of cup products, we get $u_{1}\smile\dots\smile u_{k}=0\in H^{*}(E_{B,G};R)$, contradicting the hypothesis (2). ∎
When $B$ is a point, we can use the above theorem to get a cohomological lower bound for $\mathrm{TC}_{G}(E)$.
The product inequality for parametrized topological complexity was proved in [8, Proposition 6.1]. We now establish the corresponding equivariant analogue.
Theorem 3.7.
Let $p_{1}\colon E_{1}\to B_{1}$ be a $G_{1}$-fibration and $p_{2}\colon E_{2}\to B_{2}$ be a $G_{2}$-fibration. If $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is $(G_{1}\times G_{2})$-completely normal, then
$\mathrm{TC}_{G_{1}\times G_{2}}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]\leq\mathrm{TC}_{G_{1}}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}_{G_{2}}[p_{2}\colon E_{2}\to B_{2}]-1,$
where $G_{i}$ acts on $E_{i}\times E_{i}$ diagonally for $i=1,2$; and $G_{1}\times G_{2}$ acts on $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ componentwise.
Proof.
Let $\Pi_{1}\colon(E_{1})^{I}_{B_{1}}\to E_{1}\times_{B_{1}}E_{1}$ and $\Pi_{2}\colon(E_{2})^{I}_{B_{2}}\to E_{2}\times_{B_{2}}E_{2}$ be the equivariant parametrized fibrations corresponding to $p_{1}$ and $p_{2}$, respectively. If $E:=E_{1}\times E_{2}$, $B:=B_{1}\times B_{2}$ and $p:=p_{1}\times p_{2}$ is the product $(G_{1}\times G_{2})$-fibration, then it easily checked that
$E^{I}_{B}=(E_{1})^{I}_{B_{1}}\times(E_{2})^{I}_{B_{2}}\quad\text{and}\quad E\times_{B}E=(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2})$
and the $(G_{1}\times G_{2})$-equivariant parametrized fibration $\Pi\colon E^{I}_{B}\to E\times_{B}E$ corresponding to $p$ is equivalent to the product $(G_{1}\times G_{2})$-fibration
$\Pi_{1}\times\Pi_{2}\colon(E_{1})^{I}_{B_{1}}\times(E_{2})^{I}_{B_{2}}\to(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2}).$
Since a subspace of a $(G_{1}\times G_{2})$-completely normal space is itself $(G_{1}\times G_{2})$-completely normal, it follows that $(E_{1}\times_{B_{1}}E_{1})\times(E_{2}\times_{B_{2}}E_{2})$ is $(G_{1}\times G_{2})$-completely normal. Hence,
$\mathrm{TC}_{G_{1}\times G_{2}}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]$ $=\mathrm{secat}_{G_{1}\times G_{2}}(\Pi_{1}\times\Pi_{2})$
$\leq\mathrm{secat}_{G_{1}}(\Pi_{1})+\mathrm{secat}_{G_{2}}(\Pi_{2})-1$
$=\mathrm{TC}_{G_{1}}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}_{G_{2}}[p_{2}\colon E_{2}\to B_{2}]-1,$
by [1, Proposition 3.7]. ∎
Corollary 3.8.
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
$\mathrm{TC}_{G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]\leq\mathrm{TC}_{G}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}_{G}[p_{2}\colon E_{2}\to B_{2}]-1.$
Proof.
In Proposition 2.14, we showed that $p_{1}\times p_{2}\colon E_{2}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration if $G$ is compact Hausdorff.
If $E_{1}$ and $E_{2}$ are Hausdorff, and $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is completely normal, then [10, Lemma 3.12] implies that $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ is $(G\times G)$-completely normal. Hence, the desired inequality
$\mathrm{TC}_{G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]$ $\leq\mathrm{TC}_{G\times G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]$
$\leq\mathrm{TC}_{G}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}_{G}[p_{2}\colon E_{2}\to B_{2}]-1$
follows from Corollary 3.5 (2) and Theorem 3.7. ∎
Corollary 3.9.
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration for $i=1,2$. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
$\mathrm{TC}_{G}[p\colon E_{1}\times_{B}E_{2}\to B]\leq\mathrm{TC}_{G}[p_{1}\colon E_{1}\to B]+\mathrm{TC}_{G}[p_{2}\colon E_{2}\to B]-1,$
where $G$ acts on $E_{i}\times E_{i}$ diagonally for $i=1,2$; and $G\times G$ acts on $(E_{1}\times E_{1})\times(E_{2}\times E_{2})$ componentwise.
Proof.
Note the following diagram
[…]
is a pullback in the category of $G$-spaces, where $\Delta\colon B\to B\times B$ is the diagonal map. In Corollary 2.15, we showed that $p$ is a $G$-fibration. Hence, the desired inequality
$\mathrm{TC}_{G}[p\colon E_{1}\times_{B}E_{2}\to B]$ $\leq\mathrm{TC}_{G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B\times B]$
$\leq\mathrm{TC}_{G}[p_{1}\colon E_{1}\to B]+\mathrm{TC}_{G}[p_{2}\colon E_{2}\to B]-1$
follows from [12, Proposition 4.6] and Corollary 3.8. ∎
4. Invariant Parametrized Topological Complexity
In this section, we introduce the main object of our study, the invariant parametrized topological complexity.
Suppose $p\colon E\to B$ is a $G$-fibration. Define the space
$E^{I}_{B}\times_{E/G}E^{I}_{B}:=\{(\gamma,\delta)\in E^{I}_{B}\times E^{I}_{B}\mid G\cdot\gamma(1)=G\cdot\delta(0)\}.$
That is the following diagram
[…]
is a pullback. For each path $\alpha\in E^{I}_{B}$, let $b_{\alpha}$ denote the element in $B$ such that $\alpha$ take values in the fibre $p^{-1}(b_{\alpha})$. Define the map
$\Psi\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to E\times_{B/G}E,\quad\text{by}\quad\Psi(\gamma,\delta)=(\gamma(0),\delta(1)).$ (6)
The map $\Psi$ is well-defined as $\gamma(1)=g\cdot\delta(0)$ for some $g\in G$ and $\gamma,\delta\in E^{I}_{B}$ implies that $b_{\gamma}=g\cdot b_{\delta}$. Hence, $p(\gamma(0))=b_{\gamma}=g\cdot b_{\delta}=g\cdot p(\delta(1))$ implies $(\gamma(0),\delta(1))\in E\times_{B/G}E$.
As $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$ are $(G\times G)$-invariant subsets of $E^{I}_{B}\times E^{I}_{B}$ and $E\times E$ respectively, we get $(G\times G)$-action on $E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $E\times_{B/G}E$, and $\Psi$ becomes a $(G\times G)$-equivariant map.
Proposition 4.1.
If $p\colon E\to B$ is a $G$-fibration, then the map $\Psi\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to E\times_{B/G}E$ is a $(G\times G)$-fibration.
Proof.
Suppose $E^{I}_{B}\to E\times_{B}E$ is the equivariant parametrized fibration corresponding to $p$. Suppose $\widehat{p}\colon E^{I}_{B}\times E^{I}_{B}\to(E\times_{B}E)\times(E\times_{B}E)$ is the product $(G\times G)$-fibration. Define
$S:=\{(e_{1},e_{2},e_{3},e_{4})\in(E\times_{B}E)\times(E\times_{B}E)\mid(\gamma(1),\delta(0))\in E\times_{E/G}E\}.$
It is readily checked that $(\gamma,\delta)\in E^{I}_{B}\times_{E/G}E^{I}_{B}$ if and only if $(\gamma,\delta)\in(\widehat{p})^{-1}(S)$. Since $S$ is $(G\times G)$-invariant, it follows that the restriction
$\left.\widehat{p}\right|_{E^{I}_{B}\times_{E/G}E^{I}_{B}}\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to S$
is a $(G\times G)$-fibration.
Now consider the pullback diagram
${E\times_{B}E}$${E}$${E}$${B.}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{1}}$$\scriptstyle{p}$$\scriptstyle{p}$
As $p$ is a $G$-fibration, it follows that $\pi_{1}$ and $\pi_{2}$ are $G$-fibrations. Hence, the projection map $\pi_{1,4}:=\pi_{1}\times\pi_{4}\colon(E\times_{B}E)\times(E\times_{B}E)\to E\times E$, given by $(e_{1},e_{2},e_{3},e_{4})\mapsto(e_{1},e_{4})$, is a $(G\times G)$-fibration. It is readily checked that $(e_{1},e_{2},e_{3},e_{4})\in S$ if and only if $(e_{1},e_{2},e_{3},e_{4})\in(\pi_{1,4})^{-1}(E\times_{B/G}E)$. Since $E\times_{B/G}E$ is $(G\times G)$-invariant, it follows that
$\left.\pi_{1,4}\right|_{S}\colon S\to E\times_{B/G}E$
is a $(G\times G)$-fibration. Hence, $\Psi=\left.\pi_{1,4}\right|_{S}\circ\left.\widehat{p}\right|_{E^{I}_{B}\times_{E/G}E^{I}_{B}}$ is a $(G\times G)$-fibration. ∎
We now introduce the main object of our study, which is a parametrized analogue of invariant topological complexity introduced by Lubawski and Marzantowicz in [25].
Definition 4.2.
Suppose $p\colon E\to B$ is a $G$-fibration. The invariant parametrized topological complexity, denoted by $\mathrm{TC}^{G}[p\colon E\to B]$ is defined as
$\mathrm{TC}^{G}[p\colon E\rightarrow B]:=\mathrm{secat}_{G\times G}\left(\Psi\right).$
The $G$-homotopy equivalence of the invariant topological complexity was established by Lubawski and Marzantowicz in [25, Proposition 2.4 and Lemma 3.8]. We will now establish the corresponding parametrized analogue. In particular, we establish the equivariant fibrewise homotopy equivalence of invariant parametrized topological complexity. We refer the reader to [12, Section 4.1] for basic information about fibrewise equivariant homotopy equivalence.
Theorem 4.3.
If $p\colon E\to B$ and $p^{\prime}\colon E^{\prime}\to B$ are $G$-fibrations which are fibrewise $G$-homotopy equivalent, then
$\mathrm{TC}^{G}[p\colon E\to B]=\mathrm{TC}^{G}[p^{\prime}\colon E^{\prime}\to B].$
Proof.
Suppose we have a fibrewise $G$-homotopy equivalence given by the following commutative diagram:
[…]
Suppose $\tilde{f}=f\times f$, $\tilde{f}^{I}(\gamma,\delta)=(f\circ\gamma,f\circ\delta)$ and $\tilde{f^{\prime}}$, $\tilde{f^{\prime}}^{I}$ are defined similarly. Note that $\tilde{f}$ and $\tilde{f^{\prime}}$ are $(G\times G)$-maps. Then we have the following commutative diagram.
[…]
Since the maps $f^{\prime}\circ f$ and $\mathrm{id}_{E}$ are fibrewise $G$-homotopy equivalent, it follows that the maps $\tilde{f^{\prime}}\circ\tilde{f}$ and $\mathrm{id}_{E\times_{B/G}E}$ are $(G\times G)$-homotopy equivalent. Then, using [12, Lemma 4.10(2)], we obtain the inequality
$\mathrm{TC}^{G}[p\colon E\to B]=\mathrm{secat}_{G\times G}(\Psi)\leq\mathrm{secat}_{G\times G}(\Psi^{\prime})=\mathrm{TC}^{G}[p^{\prime}\colon E^{\prime}\to B].$
Similarly, we can derive the reverse inequality, which completes the proof. ∎
The next proposition shows that the invariant parametrized topological complexity of a $G$-fibration is a generalization of both the parametrized topological complexity of a fibration [8] and the invariant topological complexity of a $G$-space [25].
Proposition 4.4.
Suppose $p\colon E\to B$ is a $G$-fibration.
  1. (1)
    If $G$ acts trivially on $E$ and $B$, then $\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[p\colon E\rightarrow B]$.
  2. (2)
    If $B=\{*\}$, then $\mathrm{TC}^{G}[p\colon E\rightarrow\{*\}]=\mathrm{TC}^{G}(E)$.
Proof.
(1) If $G$ acts trivially on $E$, then $\pi_{E}\colon E\to E/G$ is the identity map. Hence, $E^{I}_{B}\times_{E/G}E^{I}_{B}=E^{I}_{B}\times_{E}E^{I}_{B}$ which is homeomorphic to $E^{I}_{B}$ via the map $(\gamma,\delta)\mapsto\gamma*\delta$, where $\gamma*\delta$ is the concatenation of paths $\gamma$ and $\delta$. The inverse of this homeomorphism is given by $\alpha\mapsto\left(\left.\alpha\right|_{\left[0,\frac{1}{2}\right]},\left.\alpha\right|_{\left[\frac{1}{2},1\right]}\right)$ for $\alpha\in E^{I}_{B}$. If $G$ acts trivially on $B$, then $\pi_{B}\colon B\to B/G$ is the identity map. Hence, $E\times_{B/G}E=E\times_{B}E$. Therefore, the fibration $\Psi$ is given by
$\Psi\colon E^{I}_{B}\to E\times_{B}E,\quad\Psi(\alpha)=(\alpha(0),\alpha(1)).$
Hence, we get $\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[p\colon E\rightarrow B]$.
(2) If $B=\{*\}$, then $E^{I}_{B}=E^{I}$ and $E\times_{B/G}E=E\times E$. Hence, the fibration $\Psi$ is given by
$\Psi\colon E^{I}\times_{E/G}E^{I}\to E\times E,\quad\Psi(\gamma,\delta)=(\gamma(0),\delta(1)).$
Therefore, $\mathrm{TC}^{G}[p\colon E\rightarrow\{*\}]=\mathrm{TC}^{G}(E)$. ∎
Proposition 4.5.
Let $p\colon B\times F\to B$ be the trivial $G$-fibration with $G$ acting trivially on $F$. Then
$\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}(F).$
Proof.
Let $E=B\times F$. Then note that $E^{I}_{B}=B\times F^{I}$ and $E\times_{B/G}E=(B\times_{B/G}B)\times(F\times F)$. As $E/G=(B\times F)/G=(B/G)\times F$, we have
$E^{I}_{B}\times_{E/G}E^{I}_{B}=(B\times_{B/G}B)\times(F^{I}\times_{F}F^{I})\cong_{G}(B\times_{B/G}B)\times F^{I},$
where the last $G$-homeomorphism is induced by $(\gamma,\delta)\in F^{I}\times_{F}F^{I}\mapsto\gamma*\delta\in F^{I}$. Then it follows that the fibration $\Psi$ corresponding to $p$ is given by $\Psi=\mathrm{id}_{B\times_{B/G}B}\times e_{F}$, where $e_{F}\colon F^{I}\to F\times F$ is the free path space fibration corresponding to $F$. Thus, we obtain
$\mathrm{TC}^{G}[p\colon E\to B]=\mathrm{secat}_{G\times G}(\Psi)=\mathrm{secat}_{G\times G}(\mathrm{id}\times e_{F})=\mathrm{secat}(e_{F})=\mathrm{TC}(F),$
since $G$ acts trivially on $F$. ∎
Remark 4.6.
In general, if $G$ acts non-trivially on $F$, then the equality
$\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}^{G}(F)$
may not hold. For example, let $E=S^{1}\times S^{1}$ and $B=S^{1}$. If $G=S^{1}$ acts on $B$ by left multiplication and diagonally on $E$, then
$\mathrm{TC}^{S^{1}}[p\colon S^{1}\times S^{1}\to S^{1}]$ $=\mathrm{TC}[p/S^{1}\colon(S^{1}\times S^{1})/S^{1}\to S^{1}/S^{1}]$ by Theorem 4.29
$=\mathrm{TC}(S^{1}\to\{*\})$
$=\mathrm{TC}(S^{1})$ by Proposition 4.4
$=2.$
But $\mathrm{TC}^{S^{1}}(S^{1})=\mathrm{TC}(\{*\})=1$ by [25, Theorem 3.10].
Suppose $\mathbb{k}(E)$ is the saturation of the diagonal $\Delta(E)$ with respect to the $(G\times G)$-action on $E\times E$, i.e.,
$\mathbb{k}(E):=(G\times G)\cdot\Delta(E)\subseteq E\times E.$
If $E\times_{E/G}E$ is the pullback corresponding to $\pi_{E}\colon E\to E/G$, i.e.,
$E\times_{E/G}E:=\{(e_{1},e_{2})\in E\times E\mid\pi_{E}(e_{1})=\pi_{E}(e_{2})\},$
then it is readily checked that $\mathbb{k}(E)=E\times_{E/G}E\subseteq E\times_{B/G}E$. Hence, we will use the notation $\mathbb{k}(E)$ and $E\times_{E/G}E$ interchangeably.
In the next theorem, we establish the parametrized analogue of [25, Lemma 3.8].
Theorem 4.7.
Suppose $p\colon E\to B$ is a $G$-fibration. For a $(G\times G)$-invariant (not necessarily open) subset $U$ of $E\times_{B/G}E$ the following are equivalent:
  1. (1)
    there exists a $(G\times G)$-equivariant section of $\Psi\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to E\times_{B/G}E$ over $U$.
  2. (2)
    there exists a $(G\times G)$-homotopy between the inclusion map $i_{U}\colon U\hookrightarrow E\times_{B/G}E$ and a $(G\times G)$-map $f\colon U\to E\times_{B/G}E$ which takes values in $E\times_{E/G}E$.
Proof.
$(1)\implies(2)$. Suppose $s=(s_{1},s_{2})\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ is a $(G\times G)$-equivariant section of $\Psi$. Let $H\colon(E^{I}_{B}\times_{E/G}E^{I}_{B})\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be given by
$H(\gamma,\delta,t)=(\gamma_{t}^{\prime},\delta_{t}^{\prime}),\quad\text{for }(\gamma,\delta)\in E^{I}_{B}\times_{E/G}E^{I}_{B},\text{ and }t\in I,$
where $\gamma_{t}^{\prime}(s)=\gamma(s+t(1-s))$ and $\delta_{t}^{\prime}(s)=\delta(s(1-t))$. It is clear that $\gamma_{t}^{\prime},\delta_{t}^{\prime}\in E^{I}_{B}$, and $\gamma_{t}^{\prime}(1)=\gamma(1)$ and $\delta_{t}^{\prime}(0)=\delta(0)$ for all $(\gamma,\delta)\in E^{I}_{B}\times_{E/G}E^{I}_{B}$ and for all $t\in I$. Hence, $H$ is well-defined. Clearly, $H$ is $(G\times G)$-equivariant such that $H(\gamma,\delta,0)=(\gamma,\delta)$ and $H(\gamma,\delta,1)=(c_{\gamma(1)},c_{\delta(0)})$, where $c_{e}$ is the constant path in $E$ taking the value $e\in E$. Then
$F:=\Psi\circ H\circ(s\times\mathrm{id}_{I})\colon U\times I\to E\times_{B/G}E$
is a $(G\times G)$-homotopy such that $F_{0}=\Psi\circ\mathrm{id}_{E^{I}_{B}\times_{E/G}E^{I}_{B}}\circ s=i_{U}$ and $F_{1}(u)=\Psi(H_{1}(s(u)))=((s_{1}(u))(1),(s_{2}(u))(0))$. As $s(u)=(s_{1}(u),s_{2}(u))\in E^{I}_{B}\times_{E/G}E^{I}_{B}$ for all $u\in U$, it follows $F_{1}(u)=((s_{1}(u))(1),(s_{2}(u))(0))\in E\times_{E/G}E$. Hence, $F_{1}$ is the desired $(G\times G)$-homotopy.
$(2)\implies(1)$. Suppose $H\colon U\times I\to E\times_{B/G}E$ is a $(G\times G)$-homotopy between $f$ and $i_{U}$. Let $s\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be the $(G\times G)$-map given by $s(u)=(c_{\pi_{1}(f(u))},c_{\pi_{2}(f(u))})$, where $\pi_{i}\colon E\times_{B/G}E\to E$ is the projection map onto the $i$-th factor. The map $s$ is well-defined since $f$ takes values in $E\times_{E/G}E$. By $G$-homotopy lifting property of $\Psi$, there exists a $(G\times G)$-homotopy $\widetilde{H}\colon U\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ such that the following diagram
[…]
commutes. Then $\Psi\circ\widetilde{H}_{1}=H_{1}=i_{U}$ implies $\widetilde{H}_{1}$ is a $(G\times G)$-equivariant section of $\Psi$ over $U$. ∎
Corollary 4.8.
For a $G$-fibration $p\colon E\to B$, we have
$\mathrm{TC}^{G}[p\colon E\to B]=~{}_{\mathbb{k}(E)}\mathrm{cat}_{G\times G}(E\times_{B/G}E).$
4.1. Properties and Bounds
Proposition 4.9.
Suppose $p\colon E\to B$ is a $G$-fibration and $B^{\prime}$ is a $G$-invariant subset of $B$. If $E^{\prime}=p^{-1}(B^{\prime})$ and $p^{\prime}\colon E^{\prime}\to B^{\prime}$ is the $G$-fibration obtained by restriction of $p$, then
$\mathrm{TC}^{G}[p^{\prime}\colon E^{\prime}\to B^{\prime}]\leq\mathrm{TC}^{G}[p\colon E\to B].$
In particular, if $b\in B^{G}$, then the fibre $F=p^{-1}(b)$ is a $G$-space and
$\mathrm{TC}^{G}(F)\leq\mathrm{TC}^{G}[p\colon E\to B].$
Proof.
Note that we have the following commutative diagram
[…]
where $\Psi^{\prime}$ and $\Psi$ are the fibrations corresponding to $p^{\prime}$ and $p$, respectively. We will now show that this diagram is a pullback.
Suppose $Z$ is a topological space with $(G\times G)$-maps $k=(k_{1},k_{2})\colon Z\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ and $h=(h_{1},h_{2})\colon Z\to E^{\prime}\times_{B^{\prime}/G}E^{\prime}$ such that $\Psi\circ k=h$. As $\Psi\circ k=h$, we have
$k_{1}(z)(0)=h_{1}(z)\in E^{\prime}\quad\text{and}\quad k_{2}(z)(1)=h_{2}(z)\in E^{\prime}.$
As $k(z)=(k_{1}(z),k_{2}(z))\in E^{I}_{B}\times_{E/G}E^{I}_{B}$, we have
$p(k_{1}(z)(t))=b_{k_{1}(z)},\quad p(k_{2}(z)(t))=b_{k_{2}(z)}\quad\text{and}\quad k_{1}(z)(1)=g_{k(z)}\cdot k_{2}(z)(0)$
for some $b_{k_{1}(z)},b_{k_{2}(z)}\in B$, $g_{k(z)}\in G$ and for all $t\in I$.
Note that $b_{k_{1}(z)}=p(k_{1}(z)(t))=p(k_{1}(z)(0))=p(h_{1}(z))$ implies $b_{k_{1}(z)}\in B^{\prime}$ since $h_{1}(z)\in E^{\prime}=p^{-1}(B^{\prime})$. Hence, $k_{1}(z)\in(E^{\prime})^{I}_{B^{\prime}}$ since $k_{1}(z)(t)\in p^{-1}(b_{k_{1}(z)})\subset p^{-1}(B^{\prime})=E^{\prime}$ for all $t\in I$. Similarly, $b_{k_{2}(z)}\in B^{\prime}$ and $k_{2}(z)\in(E^{\prime})^{I}_{B^{\prime}}$. Hence, $k_{1}(z)(1)=g_{k(z)}\cdot k_{2}(z)(0)$ implies $\mathrm{Im}(k)\subseteq(E^{\prime})^{I}_{B^{\prime}}\times_{E^{\prime}/G}(E^{\prime})^{I}_{B^{\prime}}$. Hence, the diagram above is a pullback. Then the required inequality
$\mathrm{TC}^{G}[p^{\prime}\colon E^{\prime}\to B^{\prime}]=\mathrm{secat}_{G\times G}(\Psi^{\prime})\leq\mathrm{secat}_{G\times G}(\Psi)=\mathrm{TC}^{G}[p\colon E\to B].$
follows from [10, Proposition 4.3]. ∎
Proposition 4.10.
Let $p\colon E\to B$ be a $G$-fibration. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-space and
$\mathrm{cat}_{G}(F)\leq\mathrm{TC}^{G}(F)\leq\mathrm{TC}^{G}[p\colon E\to B].$
Furthermore,
  1. (1)
    if $E\times_{B/G}E$ is $(G\times G)$-connected, then
    $\mathrm{TC}^{G}[p\colon E\to B]\leq\mathrm{cat}_{G\times G}(E\times_{B/G}E).$
  2. (2)
    if $E\times_{B/G}E$ is a connected $(G\times G)$-$\mathrm{CW}$-complex, then
    $\mathrm{cat}_{G\times G}(E\times_{B/G}E)\leq\mathrm{dim}\left(\frac{E\times_{B/G}E}{G\times G}\right)+1.$
Consequently, if $E\times_{B/G}E$ is $(G\times G)$-connected $(G\times G)$-$\mathrm{CW}$-complex, then
$\mathrm{TC}^{G}[p\colon E\to B]\leq\mathrm{dim}\left(\frac{E\times_{B/G}E}{G\times G}\right)+1.$
Proof.
If $e\in E^{G}$, then $b=p(e)\in B^{G}$. Hence, by Proposition 4.9, $F:=p^{-1}(b)$ admits a $G$-action and $\mathrm{TC}^{G}(F)\leq\mathrm{TC}^{G}[p\colon E\to B].$ Observe that $e\in F^{G}$. Therefore, the inequality $\mathrm{cat}_{G}(F)\leq\mathrm{TC}^{G}(F)$ follows from [5, Proposition 2.7].
(1) Note that if $c_{e}$ is the constant path in $E$ which takes the value $e$, then $(c_{e},c_{e})\in(E^{I}_{B}\times_{E/G}E^{I}_{B})^{(G\times G)}$. Moreover, since $E\times_{B/G}E$ is $(G\times G)$-connected, it follows that
$\mathrm{TC}^{G}[p\colon E\to B]=\mathrm{secat}_{G\times G}(\Psi)\leq\mathrm{cat}_{G\times G}(E\times_{B/G}E).$
by [10, Proposition 4.4].
(2) Since $E\times_{B/G}E$ is connected and $(e,e)\in(E\times_{B/G}E)^{(G\times G)}$, it follows that
$\mathrm{cat}_{G\times G}(E\times_{B/G}E)\leq\mathrm{dim}\left(\frac{E\times_{B}E}{G\times G}\right)+1,$
by [27, Corollary 1.12].
Now the last inequality follows from $(1)$ and $(2)$. ∎
Corollary 4.11.
Let $p\colon E\to B$ be a $G$-fibration such that $\mathrm{TC}^{G}[p\colon E\to B]=1$. If $e\in E^{G}$, then the fibre $F=p^{-1}(p(e))$ is a $G$-contractible space.
Proof.
By Proposition 4.10, we have $\mathrm{cat}_{G}(F)=1$, i.e., $F$ is $G$-contractible. ∎
We now establish sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]$ to be $1$. This serves as a converse of Corollary 4.11.
Theorem 4.12.
Suppose $p\colon E\to B$ is a $G$-fibration such that $E\times_{B/G}E$ is a $G$-$\mathrm{CW}$-complex. Let $e\in E^{G}$. If the fibre $F=p^{-1}(p(e))$ satisfies either
  • $F$ is $G$-connected, $G$-contractible and $F^{G}=\{e\}$, or
  • $F$ strong $G$-deformation retracts to the point $e$,
then $\mathrm{TC}^{G}[p\colon E\to B]=1$.
Proof.
Note that
$\Psi^{-1}(e,e)=\{(\alpha,\beta)\in E^{I}_{B}\times E^{I}_{B}\mid\alpha(0)=\beta(1)=e,\alpha(1)=g\cdot\beta(0)\text{ for some }g\in G\}.$
Since $\alpha(0)=\beta(1)=e$ and $\alpha,\beta\in E^{I}_{B}$, it follows that the fibre $\Psi^{-1}(e,e)$ is $(G\times G)$-homeomorphic to
=$\mathcal{F}=\{\gamma\in F^{I}\mid\gamma(1/2)=e,\gamma(0)=g\cdot\gamma(1)\},$
where $(G\times G)$-action on $\mathcal{F}$ is given by
$((g_{1},g_{2})\cdot\gamma)(t)=\begin{cases}g_{1}\cdot\gamma(t)&0\leq t\leq 1/2,\\ g_{2}\cdot\gamma(t)&1/2\leq t\leq 1.\end{cases}$
This action is well-defined since $\gamma(1/2)=e\in E^{G}$.
Suppose $F$ is $G$-connected, $G$-contractible and $F^{G}=\{e\}$. Since $F$ is $G$-connected, we have ${}_{\{e\}}\mathrm{cat}_{G}(F)=\mathrm{cat}_{G}(F)$, see [25, Remark 2.3] and [10, Lemma 3.14]. Hence, ${}_{\{e\}}\mathrm{cat}_{G}(F)=1$ as $F$ is $G$-contractible. Thus, there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ for all $f\in F$. Let $K\colon F^{I}\times I\to F^{I}$ be the homotopy given by $K(\delta,t)(s)=H(\delta(s),t)$ for all $s,t\in I$ and $\delta\in F^{I}$. Note that $K$ is a $G$-homotopy. If $\gamma\in\mathcal{F}$, then
$g\cdot K(\gamma,t)(1/2)=g\cdot H(\gamma(1/2),t)=g\cdot H(e,t)=H(g\cdot e,t)=H(e,t)$
for all $g\in G$, i.e., $K(\gamma,t)(1/2)\in F^{G}$. Since $F^{G}=\{e\}$, we get $K(\gamma,t)(1/2)=e$ for all $t\in I$.
Suppose $F$ strong $G$-deformation retracts to the point $e$, then there exists a $G$-homotopy $H\colon F\times I\to F$ such that $H(f,0)=f$ and $H(f,1)=e$ and $H(e,t)=e$ for all $f\in F$ and $t\in I$. Then the homotopy $K$ defined on $F^{I}$ like above satisfies $K(\gamma,t)(1/2)=e$ due to the condition $H(e,t)=e$ for all $t\in I$.
Moreover,
$K(\gamma,t)(0)=H(\gamma(0),t)=H(g\cdot\gamma(1),t)=g\cdot H(\gamma(1),t)=g\cdot K(\gamma,t)(1)$
where $\gamma(0)=g\cdot\gamma(1)$. Hence, if $\gamma\in\mathcal{F}$, we have $K(\gamma,t)\in\mathcal{F}$.
Hence, in both cases, $K$ restricts to a $(G\times G)$-homotopy on $K\colon\mathcal{F}\times I\to\mathcal{F}$ such that $K(\gamma,0)=\gamma$ and $K(\gamma,1)=c_{e}$, where $c_{e}$ is the constant path in $E$ taking the value $e$. In particular, $\mathcal{F}$ is $(G\times G)$-contractible. Hence, by equivariant obstruction theory, $\Psi$ admits a $(G\times G)$-section. ∎
Later, we will also provide sufficient conditions for $\mathrm{TC}^{G}[p\colon E\to B]=1$ and its converse when the group action on the base is free, as stated in Corollary 4.31 and Corollary 4.15, respectively.
Proposition 4.13.
Suppose $p\colon E\to B$ is a $G$-fibration such that $G$ acts freely on $B$. If $K$ is a subgroup of $G$ such that $p\colon E\to B$ is also a $K$-fibration, then
$\mathrm{TC}^{K}[p\colon E\to B]\leq\mathrm{TC}^{G}[p\colon E\to B].$
Proof.
Suppose $\Psi_{K}\colon E^{I}_{B}\times_{E/K}E^{I}_{B}\to E\times_{B/K}E$ is the invariant parametrized fibration corresponding to $K$-fibration $p$. Then the following diagram
[…]
is commutative. Suppose $U$ is a $(G\times G)$-invariant open subset of $E\times_{B/G}E$ with a $(G\times G)$-equivariant section $s\colon U\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ of $\Psi$.
Define $V:=U\cap\left(E\times_{B/K}E\right)$. Then $V$ is $(K\times K)$-invariant open subset of $E\times_{B/K}E$. Suppose $(e_{1},e_{2})\in V$ and $s(e_{1},e_{2})=(\gamma,\delta)\in E^{I}_{B}\times_{E/G}E^{I}_{B}$. We claim that $s(e_{1},e_{2})=(\gamma,\delta)$ lies in $E^{I}_{B}\times_{E/K}E^{I}_{B}$. Note that $p(e_{1})=k\cdot p(e_{2})$ for some $k\in K$, as $(e_{1},e_{2})\in E\times_{B/K}E$. Since $s$ is a section of $\Psi$, we have
$b_{\gamma}=p(\gamma(0))=p(e_{1})=k\cdot p(e_{2})=k\cdot p(\delta(1))=k\cdot b_{\delta},$
where $\gamma(t)\in p^{-1}(b_{\gamma})$ and $\delta(t)\in p^{-1}(b_{\delta})$ for some $b_{\gamma},b_{\delta}\in B$ and for all $t\in I$. Since $(\gamma,\delta)\in E^{I}_{B}\times_{E/G}E^{I}_{B}$, we have $\gamma(1)=g\cdot\delta(0)$ for some $g\in G$. Hence,
$b_{\gamma}=p(\gamma(1))=p(g\cdot\delta(0))=g\cdot p(\delta(0))=g\cdot b_{\delta}.$
Thus, we get $g\cdot b_{\delta}=k\cdot b_{\delta}$. It follows that $g=k$ since $G$ acts freely on $B$. Thus, $\gamma(1)=k\cdot\delta(0)$ implies $(\gamma,\delta)\in E^{I}_{B}\times_{E/K}E^{I}_{B}$. Hence, the restriction $\left.s\right|_{V}\colon V\to E^{I}_{B}\times_{E/K}E^{I}_{B}$ is a $(K\times K)$-equivariant section of $\Psi_{K}$. ∎
Corollary 4.14.
Suppose $G$ is a compact Hausdorff topological group. If $p\colon E\to B$ is a $G$-fibration such that $G$ acts freely on $B$, then
$\mathrm{TC}^{K}[p\colon E\to B]\leq\mathrm{TC}^{G}[p\colon E\to B]$
for all closed subgroups $K$ of $G$. In particular,
$\mathrm{TC}(F)\leq\mathrm{TC}[p\colon E\to B]\leq\mathrm{TC}^{G}[p\colon E\to B],$
where $F$ is the fibre of $p$.
Proof.
Note that, by [20, Theorem 3], the map $p\colon E\to B$ is a $K$-fibration. Hence, the result follows from [8, Page 235] and Proposition 4.13. ∎
Corollary 4.15.
Suppose $G$ is a compact Hausdorff topological group and $p\colon E\to B$ is a $G$-fibration with fibre $F$ such that $G$ acts freely on $B$. If $\mathrm{TC}^{G}[p\colon E\to B]=1$, then the $F$ is contractible.
Proof.
This follows from Corollary 4.14 and the fact that $\mathrm{TC}(F)=1$ if and only if $F$ is contractible. ∎
4.1.1. Cohomological Lower Bounds
Lemma 4.16.
Suppose $p\colon E\to B$ is a $G$-fibration. Then the map $c\colon E\times_{E/G}E\to E^{I}_{B}\times_{E/G}E^{I}_{B}$, given by $c(e_{1},e_{2})=(c_{e_{1}},c_{e_{2}})$ where $c_{e_{i}}$ is the constant path in $E$ taking the value $e_{i}\in E$, is a $(G\times G)$-homotopy equivalence.
Proof.
Let $f\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to E\times_{E/G}E$ be the map given by $f(\gamma,\delta)=(\gamma(1),\delta(0))$. Then $f$ is $(G\times G)$-equivariant such that $(c\circ f)(\gamma,\delta)=(c_{\gamma(1)},c_{\delta(0)})$ and $f\circ c$ is the identity map of $E\times_{E/G}E$. Let $H\colon(E^{I}_{B}\times_{E/G}E^{I}_{B})\times I\to E^{I}_{B}\times_{E/G}E^{I}_{B}$ be the homotopy given by
$H(\gamma,\delta,t)=(\gamma^{\prime}_{t},\delta^{\prime}_{t}),$
where $\gamma^{\prime}_{t}(s)=\gamma(s+t(1-s))$ and $\delta^{\prime}_{t}(s)=\delta(s(1-t))$. Then following the proof of Theorem 4.7, we see that $H$ is well-defined, $(G\times G)$-equivariant, $H(\gamma,\delta,0)=(\gamma,\delta)$, and $H(\gamma,\delta,1)=(c_{\gamma(1)},c_{\delta(0)})$. Hence, $c\circ f$ is $(G\times G)$-homotopic to the identity map of $E^{I}_{B}\times_{E/G}E^{I}_{B}$. ∎
Note that the following diagram
[…]
is commutative, where $i\colon E\times_{E/G}E\hookrightarrow E\times_{B/G}E$ is the inclusion map. In other words, $\Psi$ is a $(G\times G)$-fibrational substitute for the $(G\times G)$-map $i$.
For ease of notation in the upcoming theorem, let $G^{2}$ denote the product $G\times G$.
Theorem 4.17.
Suppose $p\colon E\to B$ is a $G$-fibration. Suppose there exists cohomology classes $u_{1},\dots,u_{k}\in\widetilde{H}^{*}_{G^{2}}(E\times_{B/G}E;R)$ (for any commutative ring R) such that
$(i^{h}_{G^{2}})^{*}(u_{1})=\cdots=(i^{h}_{G^{2}})^{*}(u_{k})=0\quad\text{and}\quad u_{1}\smile\cdots\smile u_{k}\neq 0,$
then $\mathrm{TC}^{G}[p\colon E\to B]>k$.
Proof.
Note that $\Psi\circ c=i$ implies $(c^{h}_{G^{2}})^{*}\circ(\Psi^{h}_{G^{2}})^{*}=(i^{h}_{G^{2}})^{*}$. Since $c$ is a $(G\times G)$-homotopy equivalence (see Lemma 4.16), it follows $c^{h}_{G^{2}}$ is a homotopy equivalence. Hence, $(c^{h}_{G^{2}})^{*}$ is an isomorphism. Thus, the result follows from Theorem 2.2. ∎
Remark 4.18.
We note that any $G$-map $f\colon X\to Y$ that is a non-equivariant homotopy equivalence induces an isomorphism on the level of Borel cohomology, see [28]. Hence, for Theorem 4.17, we don’t need $c$ is a $(G\times G)$-homotopy equivalence, we only require $c$ to be a $(G\times G)$-map and a non-equivariant homotopy equivalence.
Now we give a non-equivariant cohomological lower bound for the invariant parametrized topological complexity. Let $E_{B,G^{2}}:=(E\times_{B/G}E)/(G\times G)$ and let $\mathbb{k}_{G^{2}}E\subseteq E_{B,G^{2}}$ denote the image of the saturated diagonal subspace $\mathbb{k}(E)=E\times_{E/G}E\subseteq E\times_{B/G}E$ under the orbit map $\rho\colon E\times_{B/G}E\to E_{B,G^{2}}$. By using Theorem 4.7 and following the arguments in Theorem 3.6, one can establish the following theorem. The proof is left to the reader.
Theorem 4.19.
Suppose $p\colon E\to B$ is a $G$-fibration. If there exists cohomology classes $u_{1},\dots,u_{k}\in H^{*}(E_{B,G^{2}};R)$ (for any commutative ring $R$) such that
  1. (1)
    $u_{i}$ restricts to zero in $H^{*}(\mathbb{k}_{G^{2}}E;R)$ for $i=1,\dots,k$;
  2. (2)
    $u_{1}\smile\dots\smile u_{k}\neq 0$ in $H^{*}(E_{B,G^{2}};R)$,
then $\mathrm{TC}^{G}[p\colon E\to B]>k$.
4.1.2. Product Inequalities
Theorem 4.20.
Let $p_{1}\colon E_{1}\to B_{1}$ be a $G_{1}$-fibration and $p_{2}\colon E_{2}\to B_{2}$ be a $G_{2}$-fibration. If $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is $(G_{1}\times G_{1}\times G_{2}\times G_{2})$-completely normal, then
$\mathrm{TC}^{G_{1}\times G_{2}}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]\leq\mathrm{TC}^{G_{1}}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}^{G_{2}}[p_{2}\colon E_{2}\to B_{2}]-1.$
Proof.
Let $\Psi_{1}\colon(E_{1})^{I}_{B_{1}}\times_{E_{1}/G_{1}}(E_{1})^{I}_{B_{1}}\to E_{1}\times_{B_{1}/G_{1}}E_{1}$ and $\Psi_{2}\colon(E_{2})^{I}_{B_{2}}\times_{E_{2}/G_{2}}(E_{2})^{I}_{B_{2}}\to E_{2}\times_{B_{2}/G_{2}}E_{2}$ be the invariant parametrized fibrations corresponding to $p_{1}$ and $p_{2}$, respectively. If $E:=E_{1}\times E_{2}$, $B:=B_{1}\times B_{2}$, $G:=G_{1}\times G_{2}$, and $p:=p_{1}\times p_{2}$ is the product $G$-fibration, then it easily checked that
$E^{I}_{B}\times_{E/G}E^{I}_{B}=\left((E_{1})^{I}_{B_{1}}\times_{E_{1}/G_{1}}(E_{1})^{I}_{B_{1}}\right)\times\left((E_{2})^{I}_{B_{2}}\times_{E_{2}/G_{2}}(E_{2})^{I}_{B_{2}}\right),$
and
$E\times_{B/G}E=\left(E_{1}\times_{B_{1}/G_{1}}E_{1}\right)\times\left(E_{2}\times_{B_{2}/G_{2}}E_{2}\right),$
and the invariant parametrized fibration $\Psi\colon E^{I}_{B}\times_{E/G}E^{I}_{B}\to E\times_{B/G}E$ corresponding to $p$ is equivalent to the product fibration $\Psi_{1}\times\Psi_{2}$. Hence,
$\mathrm{TC}^{G_{1}\times G_{2}}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]$ $=\mathrm{secat}_{(G_{1}\times G_{2})\times(G_{1}\times G_{2})}(\Psi)$
$=\mathrm{secat}_{(G_{1}\times G_{1})\times(G_{2}\times G_{2})}(\Psi_{1}\times\Psi_{2})$
$\leq\mathrm{secat}_{G_{1}\times G_{1}}(\Psi_{1})+\mathrm{secat}_{G_{2}\times G_{2}}(\Psi_{2})-1$
$=\mathrm{TC}^{G_{1}}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}^{G_{2}}[p_{2}\colon E_{2}\to B_{2}]-1,$
by [1, Proposition 3.7]. ∎
The product inequality for invariant topological complexity was proved in [25, Theorem 3.18]. In the following corollary, we show that the cofibration hypothesis assumed in [25, Theorem 3.18] can be removed by using Theorem 4.20.
Corollary 4.21.
Suppose $X$ is a $G$-space and $Y$ is a $H$-space. If $X\times X\times Y\times Y$ is $(G\times G\times H\times H)$-completely normal, then
$\mathrm{TC}^{G\times H}(X\times Y)\leq\mathrm{TC}^{G}(X)+\mathrm{TC}^{H}(Y)-1.$
Proof.
Note that $X\to\{*_{1}\}$ is a $G$-fibration and $Y\to\{*_{2}\}$ is a $H$-fibration. Hence,
$\mathrm{TC}^{G\times H}(X\times Y)$ $=\mathrm{TC}^{G\times H}[X\times Y\to\{*_{1}\}\times\{*_{2}\}]$
$\leq\mathrm{TC}^{G}[X\to\{*_{1}\}]+\mathrm{TC}^{H}[Y\to\{*_{2}\}]-1$
$=\mathrm{TC}^{G}(X)+\mathrm{TC}^{H}(Y)-1,$
by Proposition 4.4 and Theorem 4.20. ∎
The proof of the following corollary is similar to Corollary 3.8 and can be shown using Corollary 4.14 and Theorem 4.20.
Corollary 4.22.
Suppose $p_{i}\colon E_{i}\to B_{i}$ is a $G$-fibration such that $G$ acts on $B_{i}$ freely for $i=1,2$. If $G$ is compact Hausdorff, then $p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}$ is a $G$-fibration, where $G$ acts diagonally on the spaces $E_{1}\times E_{2}$ and $B_{1}\times B_{2}$. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
$\mathrm{TC}^{G}[p_{1}\times p_{2}\colon E_{1}\times E_{2}\to B_{1}\times B_{2}]\leq\mathrm{TC}^{G}[p_{1}\colon E_{1}\to B_{1}]+\mathrm{TC}^{G}[p_{2}\colon E_{2}\to B_{2}]-1.$
The proof of the following corollary is similar to that of Corollary 3.9 and follows from Proposition 4.9 and Corollary 4.22.
Corollary 4.23.
Suppose $p_{i}\colon E_{i}\to B$ is a $G$-fibration, for $i=1,2$, such that $G$ acts on $B$ freely. Let $E_{1}\times_{B}E_{2}=\{(e_{1},e_{2})\in E_{1}\times E_{2}\mid p_{1}(e_{1})=p_{2}(e_{2})\}$ and let $p\colon E_{1}\times_{B}E_{2}\to B$ be the $G$-map given by $p(e_{1},e_{2})=p_{1}(e_{1})=p_{2}(e_{2})$, where $G$ acts on $E_{1}\times_{B}E_{2}$ diagonally. If $G$ is compact Hausdorff, then $p$ is a $G$-fibration. Furthermore, if $E_{1}$ and $E_{2}$ are Hausdorff, and $E_{1}\times E_{1}\times E_{2}\times E_{2}$ is completely normal, then
$\mathrm{TC}^{G}[p\colon E_{1}\times_{B}E_{2}\to B]\leq\mathrm{TC}^{G}[p_{1}\colon E_{1}\to B]+\mathrm{TC}^{G}[p_{2}\colon E_{2}\to B]-1.$
4.2. Some technical results
In this subsection, we establish two technical results which will help us compute the invariant parametrized topological complexity of Fadell-Neuwirth fibrations in Section 5.
Definition 4.24 ([20, Section 5]).
Suppose $p\colon E\to B$ is a $G$-map and $F$ is a $G$-space. We say that $p$ is a locally trivial $G$-fibration with fibre $F$ if for each point $b\in B$ there exists a $G$-invariant open subset $U$ containing $b$ and a $G$-equivariant homeomorphism $\phi:p^{-1}(U)\to U\times F$ such that the following diagram
${p^{-1}(U)}$${U\times F}$${U}$$\scriptstyle{\phi}$$\scriptstyle{p}$$\scriptstyle{\pi_{1}}$
commutes, where $G$ acts on $U\times F$ diagonally. The map $\phi$ is called a $G$-trivialization of $p$.
Proposition 4.25.
Suppose $p\colon E\to B$ is a locally trivial $G$-fibration with fibre $F$. If $G$ acts trivially on $F$, then the induced map $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces is locally trivial with fibre $F$.
Proof.
Suppose $\phi\colon p^{-1}(U)\to U\times F$ is a $G$-trivialization of $p$ over $U$. As the quotient map $\pi_{B}\colon B\to\overline{B}$ is open, it follows $\overline{U}:=\pi_{B}(U)$ is an open subset of $\overline{B}$. Further, $U$ is $G$-invariant implies $U$ is saturated with respect to $\pi_{B}$. Hence, the restriction $\left.\pi_{B}\right|_{U}\colon U\to\overline{U}$ is an open quotient map and so is the product map $\left(\left.\pi_{E}\right|_{U}\right)\times\mathrm{id}_{F}\colon U\times F\to\overline{U}\times F$. Hence, the induced natural map $(U\times F)/G\to\overline{U}\times F$ is a homeomorphism. If $(\overline{p^{-1}(U)}):=\pi_{E}(p^{-1}(U))$, then $(\overline{p^{-1}(U)})=(\overline{p})^{-1}(\overline{U})$ since $U$ is $G$-invariant. Similarly, $\left.\pi_{E}\right|_{p^{-1}(U)}\colon p^{-1}(U)\to(\overline{p})^{-1}(\overline{U})$ is an open quotient map, and the induced natural map $p^{-1}(U)/G\to(\overline{p})^{-1}(\overline{U})$ is a homeomorphism. Hence, the homeomorphism $\phi/G\colon p^{-1}(U)/G\to(U\times F)/G$ induced by $\phi$ gives a trivialization
$\overline{\phi}\colon(\overline{p})^{-1}(\overline{U})\to\overline{U}\times F$
for $\overline{p}$ over $\overline{U}$. As $p$ is surjective, it follows $\overline{p}$ is locally trivial with fibre $F$. ∎
As noted in Theorem 2.5, the induced map $\overline{p}\colon\overline{E}\to\overline{B}$ is a $G$-fibration when $G$ is a compact Hausdorff topological group. However, to compute the invariant parametrized topological complexity of the equivariant Fadell-Neuwirth fibration, defined in [12], we require Proposition 4.25, which says $\overline{p}$ is also locally trivial. We will introduce equivariant Fadell-Neuwirth fibration and calculate their invariant parametrized topological complexity in Section 5. Now, we present one more result which will be required in Section 5.
Proposition 4.26.
Suppose $p\colon E\to B$ is a fibre bundle with fibre $F$, where the spaces $E,B,F$ are $\mathrm{CW}$-complexes. Then
$\mathrm{hdim}(E)\leq\mathrm{hdim}(B)+\dim(F).$
Proof.
Since $p$ is locally trivial, it follows that $\dim(E)\leq\dim(B)+\dim(F)$. In particular, $\mathrm{hdim}(E)\leq\dim(B)+\dim(F)$. If $h\colon B^{\prime}\to B$ is a homotopy equivalence and $E^{\prime}$ is the pullback of $E$ along $h$, then $E^{\prime}$ is a fibre bundle over $B^{\prime}$ with fibre $F$. Thus, we have $\dim(E^{\prime})\leq\dim(B^{\prime})+\dim(F)$. Note that $E^{\prime}$ is homotopy equivalent to $E$ as $h$ is a homotopy equivalence. Hence, we get $\mathrm{hdim}(E)\leq\dim(B^{\prime})+\dim(F)$ and the result follows. ∎
4.3. Invariance Theorem
Suppose $p\colon E\to B$ is a $G$-fibration such that the induced map $\overline{p}\colon\overline{E}\to\overline{B}$ between the orbit spaces is a fibration. If $\overline{\Pi}\colon(\overline{E})^{I}_{\overline{B}}\to\overline{E}\times_{\overline{B}}\overline{E}$ is the parametrized fibration induced by $\overline{p}\colon\overline{E}\to\overline{B}$, then we have a commutative diagram
[…]
where $f(\gamma,\delta)=\overline{\gamma}*\overline{\delta}$, where $\overline{\gamma}=\pi_{E}\circ\gamma$.
Lemma 4.27.
The restriction $\pi_{E}\times\pi_{E}\colon E\times_{B/G}E\to\overline{E}\times_{\overline{B}}\overline{E}$ is an open quotient map.
Proof.
As $\pi_{E}\colon E\to\overline{E}$ is an open quotient map, it follows $\pi_{E}\times\pi_{E}\colon E\times E\to\overline{E}\times\overline{E}$ is also an open quotient map. The subset $E\times_{B/G}E$ of $E\times E$ is saturated with respect to $\pi_{E}\times\pi_{E}$, since $E\times_{B/G}E$ is $(G\times G)$-invariant. Thus, $\pi_{E}\times\pi_{E}\colon E\times_{B/G}E\to(\pi_{E}\times\pi_{E})(E\times_{B/G}E)$ is an open quotient map. Note that
$(\overline{e_{1}},\overline{e_{2}})\in\overline{E}\times_{\overline{B}}\overline{E}$ $\iff\overline{p}(\overline{e_{1}})=\overline{p}(\overline{e_{2}})\in\overline{B}$
$\iff\overline{p(e_{1})}=\overline{p(e_{2})}\in\overline{B}$
$\iff p(e_{1})=g\cdot p(e_{2})\text{ for some }g\in G$
$\iff(e_{1},e_{2})\in E\times_{B/G}E.$
Hence, the result follows. ∎
Proposition 4.28.
Suppose $p\colon E\to B$ is a $G$-fibration such that $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration. Then
$\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]\leq\mathrm{TC}^{G}[p\colon E\rightarrow B].$
Proof.
Suppose $U$ is a $(G\times G)$-invariant open subset of $E\times_{B/G}E$ with a $(G\times G)$-equivariant section $s$ of $\Psi$ over $U$. Then $\overline{U}:=(\pi_{E}\times\pi_{E})(U)$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$, by Lemma 4.27. As $U$ is $(G\times G)$-invariant, it follows $U$ is saturated with respect to $\pi_{E}\times\pi_{E}$. Hence, $\pi_{E}\times\pi_{E}\colon U\to\overline{U}$ is a quotient map. Then, by universal property of quotient maps, there exists a unique continuous map $\overline{s}\colon\overline{U}\to E^{I}_{B}$ such that the following diagram
[…]
commutes. Then
$\overline{\Pi}(\overline{s}(\overline{e_{1}},\overline{e_{2}}))=\overline{\Pi}(f(s(e_{1},e_{2})))=(\pi_{E}\times\pi_{E})(\Psi(s(e_{1},e_{2})))=(\pi_{E}\times\pi_{E})(e_{1},e_{2})=(\overline{e_{1}},\overline{e_{2}})$
implies $\overline{s}$ is a section of $\overline{\Pi}$ over $\overline{U}$. Thus, the result follows since $\pi_{E}\times\pi_{E}\colon E\times_{B/G}E\to\overline{E}\times_{\overline{B}}\overline{E}$ is surjective. ∎
Theorem 4.29.
Suppose $G$ is a compact Lie group. Let $p\colon E\to B$ be a $G$-fibration and let $\overline{p}\colon\overline{E}\to\overline{B}$ be the induced fibration between the orbit spaces. If the $G$-action on $E$ is free and $\overline{E}\times\overline{E}$ is hereditary paracompact, then
$\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}].$
Proof.
Suppose $\overline{U}$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$ with section $\overline{s}$ of $\overline{\Pi}$ over $\overline{U}$. Then, by Theorem 3.2 for the trivial group action, there exists a homotopy $\overline{H}\colon\overline{U}\times I\to\overline{E}\times_{\overline{B}}\overline{E}$ such that $\overline{H}_{0}$ is the inclusion map of $i_{\overline{U}}\colon\overline{U}\hookrightarrow\overline{E}\times_{\overline{B}}\overline{E}$ and $\overline{H}_{1}$ takes value in $\Delta(\overline{E})$.
Let $U=(\pi_{E}\times\pi_{E})^{-1}(\overline{U})$. Then $U$ is $(G\times G)$-invariant and $\overline{U}$ is hereditary paracompact. Note that the following diagram
[…]
commutes. As the $G$-action on $E$ is free, it follows the action of $G\times G$ on $E\times_{E/G}E$ (and $U$) is free. Hence, by the Covering Homotopy Theorem of Palais [7, Theorem II.7.3] and Lemma 4.27, it follows there exists a $(G\times G)$-homotopy $H\colon U\times I\to E\times_{B/G}E$ such that $H_{0}=i_{U}\colon U\hookrightarrow E\times_{B/G}E$ and $(\pi_{E}\times\pi_{E})\circ H=\overline{H}\circ((\pi_{E}\times\pi_{E})\times\mathrm{id}_{I})$. As $\overline{H}_{1}$ takes value in $\Delta(\overline{E})$, it follows $H_{1}$ takes values in $E\times_{E/G}E$. Hence, by Theorem 4.7, we get a $(G\times G)$-equivariant section of $\Psi$ over $U$. Thus, $\mathrm{TC}^{G}[p\colon E\rightarrow B]\leq\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]$. ∎
We note that the main theorem in Lubawski and Marzantowicz’s paper, as stated in Theorem 2.25, can be recovered from Theorem 4.29 by taking the base space $B$ to be a point. Now, we state some corollaries of this theorem.
Corollary 4.30.
Suppose $G$ is a compact Lie group. Let $p\colon E\to B$ be a $G$-fibration and let $\overline{p}\colon\overline{E}\to\overline{B}$ be the induced fibration between the orbit spaces. If the $G$-action on $B$ is free and $\overline{E}\times\overline{E}$ is hereditary paracompact, then
$\mathrm{TC}(F)\leq\mathrm{TC}[p\colon E\rightarrow B]\leq\mathrm{TC}^{G}[p\colon E\rightarrow B]=\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}],$
where $F$ is the fibre of $p$.
Proof.
Observe that $G$ acts freely on $E$. Hence, the result follows from Corollary 4.14 and Theorem 4.29. ∎
Corollary 4.31.
Suppose $G$ is a compact Lie group and $p$ is locally trivial $G$-fibration with fibre $F$, such that $G$ acts trivially on $F$, $G$ acts freely on $B$ and $\overline{E}\times\overline{E}$ is hereditary paracompact. If $F$ is contractible and $\overline{E}\times_{\overline{B}}\overline{E}$ is homotopy equivalent to a $\mathrm{CW}$-complex, then $\mathrm{TC}^{G}[p\colon E\to B]=1$.
Proof.
By Proposition 4.25, we have $\overline{p}\colon\overline{E}\rightarrow\overline{B}$ is a locally trivial fibration with fibre $F$. We note that the fibre of the parametrized fibration $\overline{\Pi}\colon(\overline{E})^{I}_{\overline{B}}\to\overline{E}\times_{\overline{B}}\overline{E}$ induced by $\overline{p}\colon\overline{E}\to\overline{B}$ is the loop space $\Omega F$, which is contractible since $F$ is contractible. Hence, $\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]=1$ by obstruction theory. Thus, the result follows from the Invariance Theorem 4.29. ∎
5. Computations for equivariant Fadell-Neuwirth fibrations
In this section, we provide estimates for the invariant parametrized topological complexity of equivariant Fadell-Neuwirth fibrations. The ordered configuration space of $s$ points on $\mathbb{R}^{d}$, denoted by $F(\mathbb{R}^{d},s)$, is defined as
$F(\mathbb{R}^{d},s):=\{(x_{1},\dots,x_{s})\in(\mathbb{R}^{d})^{s}\mid x_{i}\neq x_{j}~{}\text{ for }~{}i\neq j\}.$
Definition 5.1 ([15]).
The maps
$p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)\quad\text{defined by}\quad p(x_{1},\dots,x_{s},y_{1},\dots,y_{t})=(x_{1},\dots,x_{s})$
are called Fadell-Neuwirth fibrations.
The space $F(\mathbb{R}^{d},s+t)$ admits an action of the permutation group $\Sigma_{s}$ defined as follows. For $\sigma\in\Sigma_{s}$, let
$\sigma\cdot(x_{1},\dots,x_{s},y_{1},\dots,y_{t})=(x_{\sigma(1)},\dots,x_{\sigma(s)},y_{1},\dots,y_{t}).$
Similarly, $\Sigma_{s}$ acts on $F(\mathbb{R}^{d},s)$ by permuting the coordinates $(x_{1},\dots,x_{s})$. Notice that the map $p$ in Definition 5.1 is $\Sigma_{s}$-equivariant. In fact, in [12], it was demonstrated that this map is a $\Sigma_{s}$-fibration.
For the rest of the section, we will use the notation $p\colon E\to B$ for the equivariant Fadell-Neuwirth fibration. The fibre $F$ of $p$ is the configuration space of $t$ points on $\mathbb{R}^{d}\setminus\mathcal{O}_{s}$, where $\mathcal{O}_{s}$ represents the configuration of $s$ distinct points in $\mathbb{R}^{d}$. More precisely, $F=F(\mathbb{R}^{d}\setminus\mathcal{O}_{s},t)$.
Remark 5.2.
[…]
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s)\to F(\mathbb{R}^{d},s)]=\mathrm{TC}(\{*\}).$
Thus, we will assume that $t\geq 1$.
If $s=1$, then the permutation group $\Sigma_{1}$ is trivial and $F(\mathbb{R}^{d},1)=\mathbb{R}^{d}$. In particular, the fibration $p$ is trivial. Hence, it follows that
$\mathrm{TC}^{\Sigma_{1}}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ $=\mathrm{TC}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ by Proposition 4.4 (1)
$=\mathrm{TC}(F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t))$ by [8, Example 4.2]
$=\mathrm{TC}(F(\mathbb{R}^{d},t+1)),$
since $F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t)$ is homotopy equivalent to $F(\mathbb{R}^{d},t+1)$. We note that topological complexity of configuration spaces was computed by Farber and Grant in [17]. Thus, we will assume that $s\geq 2$.
The parametrized topological complexity of these fibrations were computed in [8] and [9]. In particular, they proved the following result:
Theorem 5.3 ([8, Theorem 9.1] and [9, Theorem 4.1]).
Suppose $s\geq 2$, $t\geq 1$. Then
$\mathrm{TC}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\begin{cases}2t+s,&\text{if $d$ is odd},\\ 2t+s-1&\text{if $d$ is even}.\end{cases}$
We will now demonstrate that the invariant parametrized topological complexity of the Fadell–Neuwirth fibrations coincides with that of the corresponding orbit fibrations. Furthermore, it is bounded below by the parametrized topological complexity of the Fadell–Neuwirth fibrations. This implies that the complexity of the universal motion planning algorithm is greater when the order in which obstacles (mines) are placed is irrelevant. This is something one would expect, since the motion planners, in a sense, satisfies an extra condition, i.e., they remain unchanged under the reordering of the obstacles (mines).
Theorem 5.4.
The induced map $\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}$ is a locally trivial fibration with fibre $F$. Moreover,
$\mathrm{TC}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]$ $\leq\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]$
$=\mathrm{TC}[\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}]$
Proof.
We note that the equivariant Fadell-Neuwirth fibrations are locally $\Sigma_{s}$-trivial with $\Sigma_{s}$ acting trivially on the fibre $F$, see [12, Section 5.1]. Since $F(\mathbb{R}^{d},s)$ is a manifold, it is paracompact Hausdorff. Hence, by Proposition 4.25, it follows that the induced map $\overline{p}$ is a locally trivial fibration with fibre $F$.
As the action of $\Sigma_{s}$ on $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ is free, and since $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ are manifolds, it follows that $\overline{F(\mathbb{R}^{d},s+t)}$ and $\overline{F(\mathbb{R}^{d},s)}$ are also manifolds. Thus, the result follows from Corollary 4.30. ∎
Proposition 5.5.
Suppose $p\colon E\to B$ denotes the equivariant Fadell-Neuwirth fibration with fibre $F$. Then
  1. (1)
    the space $E\times_{B/G}E$ is $(d-2)$-connected, and
  2. (2)
    $\dim\left(\overline{E}\times_{\overline{B}}\overline{E}\right)=\dim(B)+2\dim(F)=ds+2dt$.
  3. (3)
    $\mathrm{hdim}(\overline{E}\times_{\overline{B}}\overline{E})\leq\mathrm{hdim}(\overline{B})+2\dim(F)=(d-1)(s-1)+2dt$.
Proof.
(1) Since $\Sigma_{s}$ is a finite group acting freely on a Hausdorff space $B$, it follows $\pi_{B}\colon B\to B/G$ is a covering map. Hence, $\pi_{B}$ is a fibration. Thus, $\pi_{1}\colon E\times_{B/G}E\to E$ is a fibration with fibre $\coprod_{g\in G}F$ since the following diagram
${E\times_{B/G}E}$${E}$${E}$${B/G}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{B}\circ p}$$\scriptstyle{\pi_{B}\circ p}$
is a pullback. As $E$ and $F$ are $(d-2)$-connected (see discussion after the statement of Theorem 4.1 in [9]), it follows that the space $E\times_{B/G}E$ is $(d-2)$-connected.
(2) As $\overline{p}\colon\overline{E}\to\overline{B}$ is a locally trivial fibration with fibre $F$, it follows that the obvious map $\overline{E}\times_{\overline{B}}\overline{E}\to\overline{B}$ is a locally trivial fibration with fibre $F\times F$. Hence,
$\dim(\overline{E}\times_{\overline{B}}\overline{E})=\dim(F\times F)+\dim(\overline{B})=2\dim(F)+\dim(\overline{B}).$
Since $\Sigma_{s}$-action on the manifold $B$ is free, we get that $\overline{B}$ is a manifold with $\dim(\overline{B})=\dim(B)$.
(3) The homotopy dimension of the unordered configuration space $\overline{B}$ is $(d-1)(s-1)$ (see [2, Example 7.1.12] and [6, Theorem 3.13]). Hence, by Proposition 4.26, the claim follows. ∎
We would like to thank Professor Jesus Gonzalez for pointing us to an appropriate reference on the equivariant $\mathrm{CW}$-complex structure on the ordered configuration space, from which we can determine the homotopy dimension of the unordered configuration space. This result is used in the following theorem and Theorem 5.11.
We are now ready to present our computations for the invariant parametrized topological complexity of the Fadell-Neuwirth fibrations for the case $d\geq 3$.
Theorem 5.6.
Suppose $s\geq 2$, $t\geq 1$ and $d\geq 3$. Then
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\begin{cases}2t+s,&\text{if $d$ is odd},\\ \text{either }2t+s-1\text{ or }2t+s&\text{if $d$ is even}.\end{cases}$ (7)
Proof.
It suffices to show that $\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]\leq 2t+s$, since the inequality
$\mathrm{TC}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]\leq\mathrm{TC}[\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}]$
established in Theorem 5.4, together with Theorem 5.3 yields the desired lower bound.
Observe that
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\mathrm{TC}[\overline{p}\colon\overline{F(\mathbb{R}^{d},s+t)}\to\overline{F(\mathbb{R}^{d},s)}]\leq 2t+s$
if and only if $(2t+s)$-fold join
$\overline{\Pi}_{2t+s}\colon*_{2t+s}(\overline{E}^{I}_{\overline{B}})\to\overline{E}\times_{\overline{B}}\overline{E}$
of the fibration $\overline{\Pi}$ admits a global section, see [32, Theorem 3]. Note that the loop space $\Omega F$ is path-connected, since $\Omega F$ is $(d-3)$-connected and $d\geq 3$. Therefore, $*_{2t+s}\Omega F$ is simply connected, and hence $k$-simple for all $k$. Thus, the obstruction to a global section of $\overline{\Pi}_{2t+s}$ lies in the cohomology classes
$H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F)),$
see [31, Theorem 34.2]. As $*_{2t+s}\Omega F$ is $((2t+s)(d-1)-2)$-connected, we have $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for $k\leq(2t+s)(d-1)-2$.
Suppose $\mathcal{N}$ is any local coefficient system on $\overline{B}$. Let $\mathcal{M}$ be the local coefficient system on $\overline{E}\times_{\overline{B}}\overline{E}$ obtained as the pullback of $\mathcal{N}$ under the fibration
$F\times F\hookrightarrow\overline{E}\times_{\overline{B}}\overline{E}\to\overline{B}.$ (8)
Consider the Serre spectral sequence $E^{p,q}_{r}$ with local coefficients for the fibration (8), see [30, Theorem 2.9]. Then $E_{2}^{p,q}=H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))$, and the spectral sequence converges to $H^{p+q}(\overline{E}\times_{\overline{B}}\overline{E},\mathcal{M})$. If either $p>\mathrm{hdim}(\overline{B})$ or $q>\mathrm{hdim}(F\times F)$, then $H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))=0$. Hence,
$H^{p+q}(\overline{E}\times_{\overline{B}}\overline{E},\mathcal{M})=0\text{ if }p+q$ $>\mathrm{hdim}(\overline{B})+\mathrm{hdim}(F\times F)$
$=(d-1)(s-1)+2(d-1)t$
$=(2t+s-1)(d-1),$
see [9, Equation 4.1] for the homotopy dimension of $F$. Since $d\geq 3$, the space $F\times F$ is simply connected. Consequently, the induced map $\pi_{1}(\overline{E}\times_{\overline{B}}\overline{E})\to\pi_{1}(\overline{B})$ is an isomorphism. This can be deduced from the long exact sequence of homotopy groups corresponding to fibration (8). This implies that every local coefficient on $\overline{E}\times_{\overline{B}}\overline{E}$ is a pullback of a local coefficient system on $\overline{B}$. As a result, we have $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for $k>(2t+s-1)(d-1)-1=(2t+s)(d-1)-d$. Thus, the obstruction classes vanishes for all $k$. ∎
We would like to thank Professor Mark Grant for suggesting the use of obstruction theory in the above proof to improve our earlier bound, and for explicitly explaining how to apply the cohomological dimension of the unordered configuration space in the Serre spectral sequence to show that the obstructions vanish.
We now turn our attention to the case $d=2$. For $d=2$, we note that the spaces $E$ and $B$ are aspherical, see [23, Lemma 3.4]. Since the maps $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps, it follows that $\overline{E}$ and $\overline{B}$ are also aspherical, as $E\to\overline{E}$ and $B\to\overline{B}$ are covering maps. Therefore, we can apply the techniques developed by Grant in [22] to compute $\mathrm{TC}[\overline{p}\colon\overline{E}\to\overline{B}]$, as the map $\overline{p}\colon\overline{E}\to\overline{B}$ is a fibration of aspherical spaces with path-connected fibre $F$. More precisely, this is equivalent to computing the topological complexity of the group epimorphism $\overline{p}_{*}\colon\pi_{1}(\overline{E})\to\pi_{1}(\overline{B})$. First, we recall some definitions that will be useful for our discussion.
Definition 5.7 ([22, Definition 3.1]).
We say that a pointed map $f\colon X\to Y$ realizes the group homomorphism $\rho\colon G\to H$ if $X$ is a $K(G,1)$ space, $Y$ is a $K(H,1)$ space, and $f$ induces the homomorphism $\rho$ on fundamental groups.
Definition 5.8 ([13]).
Suppose $G$ is a group and $K(G,1)$ is the corresponding Eilenberg-MacLane space.
  • Then $G$ is said to be of type $F$ if $K(G,1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex.
  • Then $G$ is said to be of type $FP$ if there exists a finite length resolution of $\mathbb{Z}$ by finitely generated projective $\mathbb{Z}G$-modules.
We note that a group of type $F$ is of type $FP$, see [13, Page 171].
Definition 5.9 ([13, Theorem 3.6]).
A group $G$ is said to be a duality group of cohomological dimension $n$ if $G$ is of type $FP$, and there exists a $\mathbb{Z}G$-module $D$ such that
$H^{i}(G,\mathbb{Z}G)\cong\begin{cases}0&\text{for }i\neq n,\\ D&\text{for }i=n.\end{cases}$
If $D$ can be chosen to have underlying abelian group $\mathbb{Z}$, then $G$ is said to be a Poincaré duality group of cohomological dimension $n$.
We now use arguments presented in [24, Theorem 3] to prove the following lemma. This lemma, in a certain sense, generalizes part of [13, Theorem 4.3] from Poincaré duality groups to duality groups.
Lemma 5.10.
Suppose $1\to H\to G\to K\to 1$ is a short exact sequence of groups, where $H$ is a Poincaré duality group of cohomological dimension $h$. Then $H^{p}(K,\mathbb{Z}K)=0$ if and only if $H^{p+h}(G,\mathbb{Z}G)=0$. In particular, if $K$ is of type $FP$ and $G$ is a duality group of cohomological dimension $h+k$, then $K$ is a duality group of cohomological dimension $k$.
Proof.
Consider the Lyndon-Hochschild-Serre spectral sequence corresponding to the extension $1\to H\to G\to K\to 1$ with coefficients in $\mathbb{Z}G$. Then $E_{2}^{p,q}=H^{p}(K;H^{q}(H;\mathbb{Z}G))$, and the spectral sequence converges to $H^{p+q}(G;\mathbb{Z}G)$. As $\mathbb{Z}G$ is a free $\mathbb{Z}H$-module and $H$ is a Poincaré duality group of dimension $h$, it follows that
$H^{q}(H;\mathbb{Z}G)\cong H^{q}(H;\mathbb{Z}H)\otimes_{\mathbb{Z}H}\mathbb{Z}G\cong\begin{cases}0&\text{if }q\neq h,\\ \mathbb{Z}\otimes_{\mathbb{Z}H}\mathbb{Z}G\cong\mathbb{Z}K&\text{if }q=h,\end{cases}$
as $K$-modules. This implies
$E_{2}^{p,q}=H^{p}(K;H^{q}(H;\mathbb{Z}G))\cong\begin{cases}0&\text{if }q\neq h,\\ H^{p}(K;\mathbb{Z}K)&\text{if }q=h.\end{cases}$
Observe that all lattice points on the page $E_{2}$ that do not lie on the vertical line $q=h$ are zero. Since the differential $d_{r}$ maps $E_{r}^{p,q}$ to $E_{r}^{p+r,q-r+1}$, it follows that $H^{p}(K;H^{q}(H;\mathbb{Z}G))=E_{2}^{p,q}=E_{\infty}^{p,q}=H^{p+q}(G;\mathbb{Z}G)$. In particular, $H^{p}(K;\mathbb{Z}K)=0$ if and only if $H^{p+h}(G,\mathbb{Z}G)=0$. ∎
Theorem 5.11.
Suppose $s\geq 2$ and $t\geq 2$. Then
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{2},s+t)\to F(\mathbb{R}^{2},s)]=2t+s-1.$ (9)
Proof.
Suppose $B_{s+t}$ is the braid group on $(s+t)$-strands. Then, by [23, Lemma 3.4], the fundamental group of $F(\mathbb{R}^{d},s+t)/\Sigma_{s}$ is
$B_{s+t}^{\Sigma_{s}}:=\pi^{-1}(\Sigma_{s}\times\{1\}^{t}),$
where $\pi\colon B_{s+t}\to\Sigma_{s+t}$ is the canonical map. Moreover, by [23, Lemma 3.6], we have $B_{s+t}^{\Sigma_{s}}$ is a duality group of cohomological dimension $s+t-1$.
We note that the fibration $\overline{p}\colon\overline{F(\mathbb{R}^{2},s+t)}\to\overline{F(\mathbb{R}^{2},s)}$ realizes the group epimorphism $\rho\colon B_{s+t}^{\Sigma_{s}}\twoheadrightarrow B_{s}$ which forgets the last $t$ strands. Hence, it follows that
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{2},s+t)\to F(\mathbb{R}^{2},s)]$ $=\mathrm{TC}[\overline{p}\colon\overline{F(\mathbb{R}^{2},s+t)}\to\overline{F(\mathbb{R}^{2},s)}]$
$=\mathrm{TC}[\rho\colon B_{s+t}^{\Sigma_{s}}\twoheadrightarrow B_{s}],$
by Theorem 4.29 and [22, Proposition 3.5].
If $Z$ is the centre of $B_{s+t}$, then $Z$ is infinite cyclic and the centre of $B_{s+t}^{\Sigma_{s}}$ is $Z$ as well, see [23, Lemma 3.7]. Hence, by [22, Theorem 3.5], it follows that
$\mathrm{TC}[\rho\colon B_{s+t}^{\Sigma_{s}}\twoheadrightarrow B_{s}]\leq\mathrm{cd}\left(\frac{B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}}{\Delta(Z)}\right)+1,$
where $\mathrm{cd}$ denotes the cohomological dimension of a group.
Suppose $\overline{P}_{t,s}=\pi_{1}(F)$. Then we have an extension
[…]
corresponding to the fibration $F\hookrightarrow\overline{E}\xrightarrow{\overline{p}}\overline{B}$. Pulling back this extension by $\rho\colon B_{s+t}^{\Sigma_{s}}\to B_{s}$ we get an extension
$1\to\overline{P}_{t,s}\to B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}\to 1$
Taking the quotient of $B_{s+t}^{\Sigma_{s}}$ by $Z$ gives an extension
$1\to\overline{P}_{t,s}\to\frac{B_{s+t}^{\Sigma_{s}}\times_{B_{s}}B_{s+t}^{\Sigma_{s}}}{\Delta(Z)}\to\frac{B_{s+t}^{\Sigma_{s}}}{Z}\to 1$ (10)
We note that $\overline{P}_{t,s}$ and $P_{s+t}$ are duality groups with $\mathrm{cd}(\overline{P}_{t,s})=t$ and $\mathrm{cd}(P_{s+t})=s+t-1$, see [23, Lemma 3.6].
We will now show that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group with $\mathrm{cd}(B_{s+t}^{\Sigma_{s}}/Z)=s+t-2$. As $t\geq 2$, the inclusion $Z\hookrightarrow B_{s+t}^{\Sigma_{s}}$ splits. This can be seen geometrically by the projection $B_{s+t}^{\Sigma_{s}}\to P_{2}$ obtained by forgetting the first $s+t-2$ strands, where $P_{2}$ is the pure braid group on $2$ strands. Hence,
$B_{s+t}^{\Sigma_{s}}\simeq Z\times\left(B_{s+t}^{\Sigma_{s}}/Z\right).$
Note that the space $F(\mathbb{R}^{d},s+t)$ contains a finite $\mathrm{CW}$-complex $C$ which is a $\Sigma_{s+t}$-equivariant strong deformation retract, see [6, Theorem 3.13]. Hence, $C$ is also a $\Sigma_{s}$-equivariant strong deformation retract of $F(\mathbb{R}^{d},s+t)$. Hence, the homotopy equivalence $K(B_{s+t}^{\Sigma_{s}},1)\simeq_{h}F(\mathbb{R}^{d},s+t)/\Sigma_{s}\simeq_{h}C/\Sigma_{s}$ implies $K(B_{s+t}^{\Sigma_{s}},1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex. Therefore, the homotopy equivalence
$K(B_{s+t}^{\Sigma_{s}},1)\simeq_{h}K(Z,1)\times K(B_{s+t}^{\Sigma_{s}}/Z,1)$
implies $K(B_{s+t}^{\Sigma_{s}}/Z,1)$ is homotopy equivalent to a finite $\mathrm{CW}$-complex, i.e., $B_{s+t}^{\Sigma_{s}}/Z$ is a group of type $F$. Then Lemma 5.10 applied to the extension $1\to Z\hookrightarrow B_{s+t}^{\Sigma_{s}}\to B_{s+t}^{\Sigma_{s}}/Z\to 1$ implies that $B_{s+t}^{\Sigma_{s}}/Z$ is a duality group of cohomological dimension $s+t-2$. Hence, by [4, Theorem 3.5], it follows that the middle group in (10) has cohomological dimension $s+2t-2$. ∎
Remark 5.12.
Note that when $t=1$, the group $B_{s+1}^{\Sigma_{s}}/Z$ is not torsion free, as shown in [23, Proposition 4.2]. Consequently, every $\mathrm{CW}$-complex of type $K(B_{s+1}^{\Sigma_{s}}/Z,1)$ is infinite dimensional. Therefore, the argument used in the preceding theorem fails for the case $t=1$.
Conjecture 5.13.
$\mathrm{TC}^{\Sigma_{s}}[F(\mathbb{R}^{2},s+1)\to F(\mathbb{R}^{2},s)]\leq 3+s.$
We expect the above conjecture to be true since the motion planning problem for two submarines constraint to the unknown position of $s$ mines (where the order in which mines are placed does not matter) should be more complex than of a single submarine. Hence,
$\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\leq\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+2)}\to\overline{F(\mathbb{R}^{2},s)}]=3+s,$
5.13 is true if $s=1$, see Remark 5.2 and [17, Theorem 1].
We note that we can get $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\leq 4+s$, using [8, Proposition 7.2] and Proposition 5.5 (3). Moreover, $\mathrm{TC}[\overline{F(\mathbb{R}^{2},s+1)}\to\overline{F(\mathbb{R}^{2},s)}]\geq 1+s$, by Theorem 5.3.
6. Acknowledgement
We thank Professor Mark Grant for his valuable feedback on an earlier version of this article, which significantly improved the paper in various aspects. In particular, we are grateful to him for suggesting an approach to establish a cohomological lower bound via the cohomology of orbit spaces, and for helping us prove Theorem 5.6 without any additional assumptions. We also thank Professor Jesus Gonzalez for pointing us to an appropriate reference concerning the homotopy dimension of the unordered configuration space. The first author would like to acknowledge IISER Pune - IDeaS Scholarship and Siemens-IISER Ph.D. fellowship for economical support. The second author acknowledges the support of NBHM through grant 0204/10/(16)/2023/R&D-II/2789.